### v1

# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

Abstract.
The emergence of AI-augmented Data Processing Systems (DPSs) has introduced powerful semantic operators that extend traditional data management capabilities with LLM-based processing. However, these systems face fundamental reliability (a.k.a. trust) challenges, as LLMs can generate erroneous outputs, limiting their adoption in critical domains. Existing approaches to LLM constraints—ranging from user-defined functions to constrained decoding—are fragmented, imperative, and lack semantics-aware integration into query execution. To address this gap, we introduce Semantic Integrity Constraints (SICs), a novel declarative abstraction that extends traditional database integrity constraints to govern and optimize semantic operators within DPSs. SICs integrate seamlessly into the relational model, allowing users to specify common classes of constraints (e.g., grounding and soundness) while enabling query-aware enforcement and optimization strategies.
In this paper, we present the core design of SICs, describe their formal integration into query execution, and detail our conception of grounding constraints, a key SIC class that ensures factual consistency of generated outputs. In addition, we explore novel enforcement mechanisms, combining proactive (constrained decoding) and reactive (validation and recovery) techniques to optimize efficiency and reliability. Our work establishes SICs as a foundational framework for trustworthy, high-performance AI-augmented data processing, paving the way for future research in constraint-driven optimizations, adaptive enforcement, and enterprise-scale deployments.
1. Introduction
Large Language Models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Shankar et al., 2024a; Liu et al., 2025; Anderson et al., 2025; Lu et al., 2025). We broadly refer to these systems as AI-augmented Data Processing Systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. For these systems to be fully adopted, guardrails must be placed around semantic operators.
LLM constraints—also known as assertions or guardrails—have been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related area of work involves constrained decoding, where constraints are applied to the LLM during its generation phase to ensure the output matches a specified structure (Beurer-Kellner et al., 2023; Dong et al., 2024b; guidance, 2022; OpenAI, 2024b; Willard and Louf, 2023; Zheng et al., 2024). While these existing approaches have proven their utility in enhancing the reliability of LLM outputs, they are imperative, fragmented, and lack awareness of operator semantics, limiting their usability and efficiency.
To address these limitations, we introduce Semantic Integrity Constraints (SICs) as a unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to traditional integrity constraints, SICs make the specification of constraints for predominant use cases easy and more advanced cases possible. They are declarative, allowing for a rich set of enforcement and optimization techniques, such as constraint pushdown, constrained decoding, fused evaluation, and adaptive retries. In addition, SICs are expressed in the same language that defines data processing and integrate seamlessly into the relational model. These design decisions are in line with previous efforts that argue for statement and solution specifications on LLM-based system components (Stoica et al., 2024) as well as an (extended) relational model in AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Madden et al., 2024; Lu et al., 2025).
The design of the right interface for SICs and the optimization of their enforcement present many exciting research opportunities and challenges. Different declarative interfaces have been proposed for semantic queries, with the majority adopting the relational model (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025). Although the community has not yet converged on an exact query language, we believe that the language for declaring semantic queries and SICs must be developed in tandem to maintain consistency and tight integration. Moreover, the constraint interface must enable users to declare SICs for the majority of use cases in a simple manner that does not require a deep understanding of the underlying implementation details. Towards this goal, the community will need to conduct user surveys to ensure that design decisions are grounded in real application requirements and do not introduce leaky abstractions. Lastly, optimizing SIC enforcement will require integrating modern techniques in LLM assertions and constrained decoding with traditional database query optimization methods. To meet ever-growing performance demands, novel approaches and extensions to existing solutions must also be developed. These solutions have the potential to accelerate the adoption of AI-augmented solutions in enterprise applications.
This paper presents our vision for SICs and sketches a general SIC framework for relational data processing with unstructured data. Section 2 introduces our proposed constraint interface, followed by Section 3, a categorization of SICs that addresses the most common use cases for LLM constraints. In Section 4, we discuss the enforcement of SICs, detail various optimizations, and highlight multiple promising research directions. Section 5 focuses on grounding constraints, a class of SICs that we are implementing in VectraFlow (Lu et al., 2025), an AI-augmented data flow system developed at Brown University. Finally, we outline several interesting open problems in Section 6.
2. Constraint Interface
We propose an interface that adopts the relational model and enables users to define SICs with the same declarative language used for semantic queries. This approach aligns with traditional methods for expressing integrity constraints, where users declare the logic of a constraint without detailing its physical implementation, allowing the DPS to choose the most efficient way to enforce SICs.
The majority of AI-augmented DPSs provide data flow interfaces, where users chain operators to define a query (Liu et al., 2025; Patel et al., 2024; Anderson et al., 2025). These systems offer a pandas-like interface in Python that exposes semantic operators separately from standard operators. This logical separation prevents declaring both semantic and standard predicates or functions within the same operator. However, recent work has introduced piped data flow syntax to SQL, enhancing its flexibility and extensibility (Shute et al., 2024). This new SQL dialect provides a basis for a unified interface between semantic and traditional queries, while also offering integrity constraints as first class citizens. Even though we design this interface for relational AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025), we expect that it can extend to DPSs with other data models too (Shankar et al., 2024a; Anderson et al., 2025).
LABEL:lst:query shows an example query that incorporates semantic functionality into pipe SQL syntax. We first describe the high-level query and then discuss the SIC-specific code later in this and the following sections. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr, biological sex, and age. The query first canonicalizes the sex attribute’s value into “Male” or “Female” (line 2) because the value could have been in another representation, e.g., “M” or “F”. The SET expression is defined using a string prefixed with the letter “p”. This expression declares a prompt string literal (p-string), which specifies functions or predicates that the DPS evaluates with an LLM. Similar to Python’s formatted string literals (f-strings), p-strings can refer to other attributes in the query within curly braces. Lines 5 to 7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Because the EXTEND operator effectively invokes a semantic map to project new attributes, each projected attribute declares a name and type using the AS keyword. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 9 to 10). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the data flow (line 15). Further note that the order of operations specified in the query does not determine the execution order. For example, the DPS should execute the filter on line 9 before summarizing the patient’s medical history on line 8 to avoid excessive computation.
Semantic query with pipe SQL syntax and SICs.
Crucially, our constraint interface extends the ASSERT operator provided by pipe SQL syntax, allowing users to declare SICs anywhere in the data flow. In the example query, assert operators are used in lines 4 and 11 and onward. These operators specify predicates that can refer to all attributes and operators defined upstream. This design enables the assert operator to validate values produced by attribute-generating semantic operators (e.g., maps, aggregates) as well as the presence or absence of tuples from nonattribute-generating operators (e.g., filters, joins, top-).
Since the predicates of the declared assert operators may not be satisfied, the DPS can retry semantic operators in the case of failure. For each semantic operator, the user can specify a retry threshold, which is the maximum number of times the DPS can retry an operator with the same input. The user can also set a retry threshold that applies to all semantic operators in the data flow by default. In LABEL:lst:query, the retry threshold for the SET operator is defined in line 3, and all other operators use the default threshold.
The user can also specify a failure mode (CONTINUE, DROP, or ABORT) for each assert operator, which determines how the DPS handles the failed input tuple(s). Similar to the retry threshold, the user can set a default failure mode. In the example query, line 4 defines the CONTINUE failure mode on the assert operator, while all other assert operators use the default. If a tuple does not satisfy the assert operator’s predicate, then the CONTINUE failure mode continues to propagate it downstream, while DROP drops it. Note that CONTINUE and DROP correspond to the (soft) Suggest and (hard) Assert constructs in DSPy Assertions (Singhvi et al., 2024), respectively. Lastly, ABORT aborts the entire query if a tuple does not satisfy the predicate. While the semantics of aborting are straightforward in a transactional context, in a data flow context, this means that the system will not only drop the tuple that violated the SIC but also discard all of its dependent tuples, such as those that would be in the same aggregation window. In any case, the DPS will log the appropriate errors and warnings when a tuple fails a SIC or when tuples are dropped to notify users of failures.
3. Constraint Classes
We categorize SICs into different classes (Table 1) that are designed to capture the predominant use cases for LLM constraints:
| Class | Description | Examples |
| Domain | Value is from a specified domain | Data types, structured output, multiple choices, numerical ranges, or length constraints |
| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHRs |
| Grounding | Value is grounded in the source values | Extracted results or summaries are factual w.r.t. original EHRs |
| Soundness | LLM’s reasoning process is logically sound | Predictions of patient diagnoses are based on sound reasoning |
| Relevance | Value is relevant to the given task | Extracted histories from EHRs are medical (desired) rather than social (undesired) |
| Assertions | Arbitrary predicate holds in the data flow | All included URLs in generated research reports lead to valid websites |
Domain Constraints. In the context of SICs, domain constraints remain similar to their traditional counterpart, checking whether an attribute’s value is from a specified domain. For example, a domain constraint can enforce that a value is of a certain data type or structure, within a numerical range, from a finite set of possible values, or less than a certain length. (Type definitions of generated attributes from semantic operators implicitly specify domain constraints.) In the example query from LABEL:lst:query, the first conjunct in line 4 and the conjunct in line 14 explicitly define domain constraints, restricting the sex to either “Male” or “Female” and the summary’s length to less than 1000 characters, respectively. This class of SICs comprise the majority of constraint use cases (Liu et al., 2024). Importantly, the DPS can enforce domain constraints programmatically without relying on an LLM validator, providing correctness guarantees on their evaluation.
Inclusion and Exclusion (IE) Constraints. IE constraints determine if particular terms, content, topics, or abstract features like tones and styles are included or excluded from an attribute’s value. For instance, these constraints ensure that an attending physician’s name is present in an EHR, while avoiding the inclusion of PII. Users can specify these constraints with a set of values that indicate what must be included or excluded. This set can be derived from a fixed collection or another attribute’s value. The user can opt for checks based on exact keyword matches or semantic similarity, even using regular expressions for matches on emails or phone numbers. Users can also define IE constraints with a p-string for content, topics, tones, or styles, as shown in line 13 of the example query, which excludes test results from the medical history summary. In this case, the DPS employs an LLM to enforce the constraint, while lightweight models can be used to handle PII or toxic content checks, such as NER models, LSTMs, or feature-based classifiers.
Grounding Constraints. These constraints target both extractive and abstractive use cases, where users want to check whether the output values of semantic operators are grounded in the source values. For instance, line 11 from the example query ensures that the extracted physical exam and lab results appear verbatim in the original EHR (extractive use case). Similarly, line 12 validates that the generated summary of the patient’s medical history is factually consistent with the EHR (abstractive use case). We discuss grounding constraints in greater detail in Section 5.
Soundness Constraints. These constraints verify whether the underlying LLM’s reasoning process is logically sound in that its premises are true and its reasoning steps are valid. This class of constraints targets nonattribute-generating semantic operators and semantic transformations from one or more values to a new value. For such cases, it is nontrivial to simply evaluate predicates on the operator’s output without invoking an LLM validator to redo the task. Instead of this similarly error-prone approach, the DPS uses an LLM validator to check the underlying LLM’s reasoning process, e.g., chain-of-thought (CoT) (Wei et al., 2022). For example, consider the semantic filter (defined on lines 9 and 10), which selects tuples where the patient is predicted to have sepsis based on their medical condition. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to outline sepsis symptoms and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator to ensure that the premises are grounded in the input and that the reasoning steps are valid. This step is defined with the assert operator in line 15. Similarly, a soundness constraint is specified in the second conjunct of line 4, which ensures that sex is canonicalized correctly. If the underlying LLM already performs well in a closed-book setting and the retrieved documents do not provide enough context to produce a valid answer, another approach is to have the LLM generate its CoT first and then verify each proposed premise with the input tuple(s) or documents retrieved afterwards. The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023). This is an avenue for future research.
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given semantic operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query; e.g., the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than their social history. Relevance constraints are evaluated by passing the operator’s prompt, input, and output to an LLM validator for verification.
Assertions. Like assertions in traditional integrity constraints, these constraints enable users to specify arbitrary predicates that must hold in the data flow. The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the primitives described previously.
4. Constraint Enforcement
Once users define SICs, the DPS will identify the most efficient enforcement method. The DPS initially moves all specified SICs to the earliest evaluation point to minimize unnecessary computation, akin to predicate pushdown in logical query optimization. If a SIC involves checking attributes, the DPS will move the constraint to the earliest data flow point where all attributes are defined, usually right after a semantic map or aggregate operator, since these are the only attribute-generating operators in our extended relational model. For SICs that verify operator processing logic, like soundness constraints, the DPS positions them at their associated semantic operator. Additionally, the DPS converts specified SICs into text and appends them to the LLM’s prompt, ensuring that the semantic operator’s LLM is aware of the constraints on its results. At this point, the DPS will follow either a proactive or reactive constraint enforcement strategy.
4.1. Proactive Enforcement
The DPS can proactively enforce certain types of SICs by pushing their enforcement into the LLM’s decoding process via constrained decoding (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022). In doing so, the output of the LLM will be guaranteed to satisfy the specified constraints by construction; thus, requiring no further checks on its output. Provided the constrained decoding overhead remains low, the DPS should proactively enforce SICs where feasible to optimize performance. Concerns around distorting the LLM’s distribution can be addressed with recent methods as well (Park et al., 2024).
Grammar-Based Constrained Decoding. An effective method for using constrained decoding to ensure specific SICs is by encoding them as a formal grammar. This is similar to how libraries (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) allow users to provide a JSON Schema (Org, 2010) to which the LLM’s output adheres. These libraries convert the JSON Schema into a context-free grammar or regular expression and apply a push-down or finite automaton, respectively, to guide the output. This allows the DPS to enforce certain domain constraints. Beyond exploring more efficient grammar-based decoding methods, open challenges include determining which SICs can be expressed as grammars and how to represent and compile multiple SICs into a grammar if they are defined on the same attribute.
Direct Constrained Decoding. An alternative constrained decoding method is to enforce SICs directly, without translating them into a formal grammar. Although we do not anticipate that this will allow the DPS to enforce all SICs, we expect it to enable a broader range of constraints than grammar-based methods. Previous research (Beurer-Kellner et al., 2023) introduced a direct, more flexible approach, but is significantly slower than current grammar-based techniques (Zheng et al., 2024). Whether this more flexible enforcement method can be implemented to efficiently enforce SICs is an open problem.
4.2. Reactive Enforcement and Recovery Actions
The DPS can also reactively enforce SICs by first detecting a constraint violation and then performing a recovery action.
4.2.1. Violation Detection
To detect a SIC violation, the DPS simply evaluates its associated predicate and checks if the predicate returns false. However, if multiple SICs are specified in the data flow, there are many opportunities to optimize their collective evaluation.
Constraint Reordering. If the DPS needs to execute multiple SICs sequentially (e.g., if the user specifies the constraints on the same attribute), then it will reorder the constraints to reduce execution costs. This is similar to predicate reordering in traditional database optimization. For example, the DPS can reorder SICs based on the cost and selectivity of each SIC’s predicate (Abadi et al., 2003). The cost of a SIC can be based on a single metric, such as execution time, or even a linear combination of cost metrics. The selectivity of a SIC is the probability that an output tuple fails the associated predicate. The DPS can learn the cost and selectivity of each SIC via offline or online methods by computing lightweight statistics.
Constraint Fusion. The DPS will further improve performance by merging SICs together into a single execution unit for certain scenarios. Akin to operator fusion, the DPS will combine code-based SICs together to reduce intermediate data materialization and unnecessary memory traffic. The DPS will also do the same for LLM-based SICs, where several constraints are evaluated by a single LLM invocation. Though, as more LLM-based constraints are fused together, the accuracy of results from the LLM validator is likely to degrade due to the increase in task complexity and input length (Levy et al., 2024). As such, finding the optimal number of constraints to combine that maintains target accuracy metrics while minimizing execution costs is an open problem.
4.2.2. Violation Recovery
To recover from a SIC violation, the DPS will retry the semantic operator with the failed input tuple(s) by using prompt or model-based techniques to improve the output quality. We outline a few approaches for retrying semantic operators. These approaches can be combined in different ways, and it is interesting to explore whether the DPS can learn their optimal combination with respect to different target metrics, such as accuracy, latency, and monetary cost.
Few-Shot Prompting. A common approach for retrying a semantic operator is to append the erroneous output and corresponding feedback message to the previous prompt on each failure (Singhvi et al., 2024; Shankar et al., 2024a). The feedback message can be a static error message specified for a code-based predicate or part of the response from an LLM validator. This refinement process provides the underlying LLM negative few-shot examples of the given task for in-context learning (Brown et al., 2020).
A complementary approach is to use the top- most similar positive few-shot examples, which are input-output pairs that previously satisfied all the specified SICs for the semantic operator. Importantly, examples are retrieved on the basis of their similarity to the LLM’s current input. The DPS can use a cache with a vector index to efficiently store and retrieve these examples. For each retry, the number of retrieved examples can increase to improve the quality of the LLM’s output. We believe that this approach will be particularly useful in data flow systems, where queries are mostly static. There are several challenges in our setting, e.g., determining ’s initial value and increment size for each retry, and example cache management policies.
Chain-of-Thought Prompting. When retrying a semantic operator, the DPS can also prompt the underlying LLM to use more reasoning steps in its CoT. By effectively scaling up inference-time computation, recent work has shown that long CoT in reasoning models generally improves the accuracy of responses (OpenAI, 2024a; DeepSeek-AI, 2025). For each retry, the DPS can thus prompt the LLM to use increasingly longer CoT to generate better outputs. It is unclear how long the initial CoT should be and by how much to increase it on each retry.
Model-Based Improvements. To improve the results of semantic operators, the DPS can invoke a more powerful LLM to generate a more accurate output, or use multiple LLMs to adopt a majority voting approach, e.g., where the same input is sent to increasingly more LLMs on each retry. In simple cases, such as when the operator outputs a classification label or a numerical value, the DPS can programmatically consider the outputs of each LLM and return the most common result or their average, respectively. In more complex scenarios, such as when the operator generates a summary, the DPS can use a separate LLM to consider all the outputs and produce the final result. An open avenue for exploration is to compare the accuracy of these two methods when various cost metrics are fixed.
5. Grounding Constraints
In this section, we describe grounding constraints in more detail to illustrate the design decisions required to integrate SICs. These constraints are being implemented in VectraFlow (Lu et al., 2025), an AI-augmented data flow system that we are developing at Brown. However, we aim to keep the presentation general since the ideas proposed here are applicable to other AI-augmented DPSs as well. We first discuss preliminaries related to grounding constraints and then describe their implementation in the rest of the section.
To understand the rationale for grounding constraints, we note a key distinction: an LLM’s response comes from its internal (parametric) knowledge, external (nonparametric) knowledge, or a combination of both. In an AI-augmented DPS, this external knowledge includes input tuples to the system or tuples retrieved from stored tables, termed source tuples, with their elements called source values. These values, considered factual, are essential to grounding constraints. Across extractive and abstractive contexts, users seek to confirm if a value generated by a semantic operator is anchored in source values. Grounding constraints facilitate these verifications.
5.1. Enforcement
When a grounding constraint is specified for an attribute, the DPS recursively applies checks to all ancestor attributes in its lineage. These grounding checks are necessary because the attribute may have been derived through a series of intermediate semantic operators from the source value(s). The DPS evaluates these checks immediately after attributes are generated, ensuring that the operator’s output is grounded in its input. In the example query from LABEL:lst:query, med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Since the grounding constraint specified on line 12 for med_hist_sum recursively defines a grounding check on med_hist, the query does not need to explicitly specify a grounding constraint on med_hist, though doing so is equivalent.
The notion of grounding differs depending on whether the operator is extractive or abstractive. For extractive operators, grounded output is precisely contained within the input. For abstractive operators, grounded output is factually consistent with the input. The DPS cannot inherently identify whether an operator is extractive or abstractive; hence, users must annotate extractive semantic maps. Without such annotations, attribute-generating operators in the data flow are presumed abstractive if they produce an attribute within the grounding constraint’s lineage. In the example query, we see the EXTRACTIVE keyword included for the extractive operations on lines 5 to 7, while no annotation is present for the abstractive operation that summarizes the medical history on line 8. Evaluating grounding checks on extractive and abstractive semantic operators necessitates different check types. The DPS will automatically assess the appropriate checks based on each operator’s semantics. We will now discuss these grounding check types.
5.2. Types of Grounding Checks
We propose three types of grounding checks with different verification mechanisms, targeting either extractive or abstractive tasks.
Exact Match. A grounding check based on exact match verifies whether an output attribute’s value is exactly contained within the input value. These checks are designed for extractive semantic map operations, where users require exact guarantees on output values, such as phys_exam, lab_res, and med_hist in the example query. The DPS can further optimize exact match verification by building inverted indices over stored relations. However, in data flow systems, evaluating these checks on values derived from input data streams are not amenable to such optimizations.
Factual Consistency Score. These checks evaluate whether an output value is grounded in the input value(s) by computing factual consistency scores with a specialized model. Each sentence in an output value is separately given with the input value(s) to the factual consistency model, and the resulting score determines whether the sentence is grounded. While these models are stochastic, they can validate summaries from abstractive semantic maps (e.g., med_hist_sum) and aggregates. Models like MiniCheck have also been shown to be more efficient than their LLM counterparts and reach GPT-4 accuracy (Tang et al., 2024).
LLM Validator. LLM-based grounding checks also target abstractive use cases, invoking an LLM validator to determine grounding. The LLM validator receives the output value and input value(s), and it returns a factual consistency score or a boolean value. Additionally, the LLM validator provides a feedback message, which can help the semantic operator output more grounded responses when retried. In fact, the DPS can combine the LLM-based check with the score-based check, where it invokes the LLM validator only if the factual consistency model classifies the output as ungrounded.
6. Open Problems
In this section, we present additional open problems that pave the way for promising research opportunities.
Adaptive Enforcement. A DPS can strategically manage performance and risk by dynamically applying SICs. By tracking SIC violations for each semantic operator, it determines when checks are necessary, thus reducing overhead. Violations incur a (normalized) penalty score reflecting their risk level, and these scores are combined into a sum. If this sum exceeds a set threshold, a retry is initiated to maintain critical integrity. This approach allows the system to adapt on the basis of each operator’s reliability, concentrating on critical validations while minimizing disruptions when operator confidence is high. Another approach involves adaptively enforcing SICs based on target metrics (e.g., accuracy, F1 score).
LLM Input Verification. While our discussion of SICs has mainly emphasized validating semantic operator outputs, it is equally crucial to validate the inputs, which may come from external sources (including a stored table). This step is necessary to avoid security risks, such as (i) prompt attacks (AI, 2024; Llama, 2024; Microsoft, 2024; Rebedea et al., 2023) or (ii) prevent unauthorized access to dark data (IBM, 2023), which refers to unprotected data hidden in a vector database that should not be used because it can pose security, privacy, or compliance risks due to its sensitive contents. An open challenge here is developing robust and scalable mechanisms to perform real-time, context-aware input verification that can effectively identify and filter malicious or unauthorized data before it reaches the LLM.
Enterprise-Wide Automated Constraint Verification. To enable a “check all” feature for semantic queries, each semantic operator in the query path must be automatically verified. Strategically determining what checks to use, where and how to place them involves examining data flow semantics and lineage to pinpoint where operator outputs affect subsequent results. The types of checks, which range from basic range checks to LLM validators, should match each operator’s semantics and risk of its output. Although our discussion has focused on single queries, enterprise settings often involve numerous queries that share common domain-specific constraints. Therefore, an important challenge is developing a mechanism to automatically capture and manage these recurring constraints in a constraint store and apply them consistently across all relevant semantic queries, ensuring enterprise-wide data integrity. To further enhance this process, the DPS should log all SIC violations, coupled with user feedback on incorrect validation results. This approach facilitates continuous LLM refinement through training, fine-tuning, or in-context learning, ultimately bolstering the accuracy of enterprise-wide constraint verification.
Leveraging LLM Progress. As LLM reliability and efficiency advance, they will influence SIC design and related trade-offs. Increases in LLM accuracy allow for more adaptivity in SIC enforcement. A critical question is: How do we determine when LLM accuracies warrant this shift? Additionally, as LLMs become more efficient, using advanced ensemble methods (e.g., confidence-based aggregation) to further improve reliability as well as maintaining LLM diversity to prevent shared errors become important.
7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases to ensure that modifications maintain data consistency. SICs build on top of traditional integrity constraints, improving the reliability of the semantic operators themselves by protecting against erroneous outputs from the underlying LLMs.
In ML, assertions were introduced as an abstraction to monitor and improve models, invoking UDFs over model outputs to flag possible errors (Kang et al., 2020). Recent works adopt similar ideas by further leveraging ML models and LLM validators to define guardrails on LLMs (Rebedea et al., 2023; Dong et al., 2024a; AI, 2023a, b; Index, 2023), with DSPy Assertions introducing an iterative refinement approach to improve output quality (Singhvi et al., 2024). Methods for validating an LLM’s CoT have been developed as well (Ling et al., 2023). In addition to these reactive constraint enforcement approaches, a related area of work uses constrained decoding to proactively enforce constraints (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022; Beurer-Kellner et al., 2023). SICs consolidate these various approaches into a unified and declarative abstraction designed for AI-augmented DPSs. Prior works like spade (Shankar et al., 2024b), which automatically synthesizes assertions, and EvalGen (Shankar et al., 2024c), which aligns LLM validators with human preferences, are complimentary to SICs and can be integrated to enhance usability and reliability.
Several recent systems have augmented traditional database operators with the semantic capabilities of LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025; Lu et al., 2025; Shankar et al., 2024a). So far, DocETL (Shankar et al., 2024a) is the only AI-augmented DPS to provide validation as a first class citizen. They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. We further their approach by proposing a declarative, relational interface for the easy specification of common constraint classes and their optimized enforcement.
Many methods have been developed to evaluate the factual consistency of LLM responses with respect to provided source documents. This problem is also known as document-level textual entailment. In addition to an exact keyword match approach that is only appropriate for extractive use cases, recent methods for evaluating factual consistency in abstractive tasks are largely categorized as those that use LLM validators and those that use specialized factual consistency models. The latter models are specifically trained to evaluate factual consistency, primarily in the context of summarization (Tang et al., 2024; Zha et al., 2023; Kryscinski et al., 2020; Goyal and Durrett, 2020; Laban et al., 2022). These models provide factual consistency scores and are more efficient than LLM-based methods, with some even reaching GPT-4 accuracy (Tang et al., 2024; Zha et al., 2023). Regardless of the specific method for evaluating factual consistency, the interface provided by SICs is designed to be agnostic of the underlying implementation.
8. Conclusions
SICs are a unified abstraction that extend traditional database integrity constraints to enhance the reliability of AI-augmented DPSs. They provide a declarative, relational interface to easily specify common classes of constraint while maintaining the flexibility to support more sophisticated use cases. Crucially, this declarative interface enables the DPS to apply a wide range of highly optimized implementations. By adopting this principled approach, SICs open new, grounded research avenues towards the goal of leveraging AI-augmented DPSs for critical workloads where reliability is paramount.
Acknowledgements.
We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback and contributions.References
- (1)
- Abadi et al. (2003) Daniel J. Abadi, Don Carney, Uğur Çetintemel, Mitch Cherniack, Christian Convey, Sangdon Lee, Michael Stonebraker, Nesime Tatbul, and Stan Zdonik. 2003. Aurora: a new model and architecture for data stream management. The VLDB Journal 12 (2003), 120–139. https://doi.org/10.1007/s00778-003-0095-z
- AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
- AI (2024) Guardrails AI. 2024. Detect Jailbreak. https://hub.guardrailsai.com/validator/guardrails/detect_jailbreak
- AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
- Anderson et al. (2025) Eric Anderson, Jonathan Fritz, Austin Lee, Bohou Li, Mark Lindblad, Henry Lindeman, Alex Meyer, Parth Parmar, Tanvi Ranade, Mehul A. Shah, et al. 2025. The Design of an LLM-powered Unstructured Analytics System. In CIDR.
- Astrahan et al. (1976) M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, K. P. Eswaran, J. N. Gray, P. P. Griffiths, W. F. King, R. A. Lorie, P. R. McJones, J. W. Mehl, G. R. Putzolu, I. L. Traiger, B. W. Wade, and V. Watson. 1976. System R: relational approach to database management. 1, 2 (June 1976), 97–137. https://doi.org/10.1145/320455.320457
- Beurer-Kellner et al. (2023) Luca Beurer-Kellner, Marc Fischer, and Martin Vechev. 2023. Prompting Is Programming: A Query Language for Large Language Models. Proc. ACM Program. Lang. 7, PLDI, Article 186 (June 2023), 24 pages. https://doi.org/10.1145/3591300
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Proceedings of the 34th International Conference on Neural Information Processing Systems (Vancouver, BC, Canada) (NIPS ’20). Curran Associates Inc., Red Hook, NY, USA, Article 159, 25 pages.
- Chamberlin et al. (1976) D. D. Chamberlin, M. M. Astrahan, K. P. Eswaran, P. P. Griffiths, R. A. Lorie, J. W. Mehl, P. Reisner, and B. W. Wade. 1976. SEQUEL 2: A Unified Approach to Data Definition, Manipulation, and Control. IBM Journal of Research and Development 20, 6 (1976), 560–575. https://doi.org/10.1147/rd.206.0560
- Chen et al. (2023) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W. Cohen. 2023. Program of Thoughts Prompting: Disentangling Computation from Reasoning for Numerical Reasoning Tasks. Transactions on Machine Learning Research (2023).
- Codd (1970) E. F. Codd. 1970. A relational model of data for large shared data banks. Commun. ACM 13, 6 (June 1970), 377–387. https://doi.org/10.1145/362384.362685
- DeepSeek-AI (2025) DeepSeek-AI. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948 [cs.CL] https://arxiv.org/abs/2501.12948
- Dong et al. (2024a) Yi Dong, Ronghui Mu, Gaojie Jin, Yi Qi, Jinwei Hu, Xingyu Zhao, Jie Meng, Wenjie Ruan, and Xiaowei Huang. 2024a. Building Guardrails for Large Language Models. arXiv:2402.01822 [cs.CL] https://arxiv.org/abs/2402.01822
- Dong et al. (2024b) Yixin Dong, Charlie F. Ruan, Yaxing Cai, Ruihang Lai, Ziyi Xu, Yilong Zhao, and Tianqi Chen. 2024b. XGrammar: Flexible and Efficient Structured Generation Engine for Large Language Models. arXiv:2411.15100 [cs.CL] https://arxiv.org/abs/2411.15100
- Eswaran and Chamberlin (1975) Kapali P. Eswaran and Donald D. Chamberlin. 1975. Functional specifications of a subsystem for data base integrity. In Proceedings of the 1st International Conference on Very Large Data Bases (Framingham, Massachusetts) (VLDB ’75). Association for Computing Machinery, New York, NY, USA, 48–68. https://doi.org/10.1145/1282480.1282484
- Goyal and Durrett (2020) Tanya Goyal and Greg Durrett. 2020. Evaluating Factuality in Generation with Dependency-level Entailment. In Findings of the Association for Computational Linguistics: EMNLP 2020, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 3592–3603. https://doi.org/10.18653/v1/2020.findings-emnlp.322
- guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
- Held et al. (1975) G. D. Held, M. R. Stonebraker, and E. Wong. 1975. INGRES: a relational data base system. In Proceedings of the May 19-22, 1975, National Computer Conference and Exposition (Anaheim, California) (AFIPS ’75). Association for Computing Machinery, New York, NY, USA, 409–416. https://doi.org/10.1145/1499949.1500029
- IBM (2023) IBM. 2023. What is dark data? Retrieved 2025-02-27 from https://www.ibm.com/think/topics/dark-data
- Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
- Kang et al. (2020) Daniel Kang, Deepti Raghavan, Peter Bailis, and Matei Zaharia. 2020. Model Assertions for Monitoring and Improving ML Models. In Proceedings of Machine Learning and Systems, Vol. 2. 481–496.
- Kryscinski et al. (2020) Wojciech Kryscinski, Bryan McCann, Caiming Xiong, and Richard Socher. 2020. Evaluating the Factual Consistency of Abstractive Text Summarization. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 9332–9346. https://doi.org/10.18653/v1/2020.emnlp-main.750
- Laban et al. (2022) Philippe Laban, Tobias Schnabel, Paul N. Bennett, and Marti A. Hearst. 2022. SummaC: Re-Visiting NLI-based Models for Inconsistency Detection in Summarization. Transactions of the Association for Computational Linguistics 10 (2022), 163–177. https://doi.org/10.1162/tacl_a_00453
- Levy et al. (2024) Mosh Levy, Alon Jacoby, and Yoav Goldberg. 2024. Same Task, More Tokens: the Impact of Input Length on the Reasoning Performance of Large Language Models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 15339–15353. https://doi.org/10.18653/v1/2024.acl-long.818
- Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 36407–36433. https://proceedings.neurips.cc/paper_files/paper/2023/file/72393bd47a35f5b3bee4c609e7bba733-Paper-Conference.pdf
- Liu et al. (2025) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano. 2025. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In CIDR.
- Liu et al. (2024) Michael Xieyang Liu, Frederick Liu, Alexander J. Fiannaca, Terry Koo, Lucas Dixon, Michael Terry, and Carrie J. Cai. 2024. ”We Need Structured Output”: Towards User-centered Constraints on Large Language Model Output. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 10, 9 pages. https://doi.org/10.1145/3613905.3650756
- Llama (2024) Meta Llama. 2024. Prompt Guard. https://www.llama.com/docs/model-cards-and-prompt-formats/prompt-guard/
- Lu et al. (2025) Duo Lu, Siming Feng, Jonathan Zhou, Franco Solleza, Malte Schwarzkopf, and Uğur Çetintemel. 2025. VectraFlow: Integrating Vectors into Stream Processing. In CIDR.
- Lyu et al. (2023) Qing Lyu, Shreya Havaldar, Adam Stein, Li Zhang, Delip Rao, Eric Wong, Marianna Apidianaki, and Chris Callison-Burch. 2023. Faithful Chain-of-Thought Reasoning. arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
- Madden et al. (2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska. 2024. Databases Unbound: Querying All of the World’s Bytes with AI. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4546–4554. https://doi.org/10.14778/3685800.3685916
- Microsoft (2024) Microsoft. 2024. Prompt Shields. https://learn.microsoft.com/en-us/azure/ai-services/content-safety/concepts/jailbreak-detection
- OpenAI (2024a) OpenAI. 2024a. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/
- OpenAI (2024b) OpenAI. 2024b. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
- Park et al. (2024) Kanghee Park, Jiayu Wang, Taylor Berg-Kirkpatrick, Nadia Polikarpova, and Loris D'Antoni. 2024. Grammar-Aligned Decoding. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 24547–24568. https://proceedings.neurips.cc/paper_files/paper/2024/file/2bdc2267c3d7d01523e2e17ac0a754f3-Paper-Conference.pdf
- Patel et al. (2024) Liana Patel, Siddharth Jha, Parth Asawa, Melissa Pan, Carlos Guestrin, and Matei Zaharia. 2024. Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data. arXiv:2407.11418 [cs.DB] https://arxiv.org/abs/2407.11418
- Rebedea et al. (2023) Traian Rebedea, Razvan Dinu, Makesh Narsimhan Sreedhar, Christopher Parisien, and Jonathan Cohen. 2023. NeMo Guardrails: A Toolkit for Controllable and Safe LLM Applications with Programmable Rails. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Yansong Feng and Els Lefever (Eds.). Association for Computational Linguistics, Singapore, 431–445. https://doi.org/10.18653/v1/2023.emnlp-demo.40
- Shankar et al. (2024a) Shreya Shankar, Tristan Chambers, Tarak Shah, Aditya G. Parameswaran, and Eugene Wu. 2024a. DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing. arXiv:2410.12189 [cs.DB] https://arxiv.org/abs/2410.12189
- Shankar et al. (2024b) Shreya Shankar, Haotian Li, Parth Asawa, Madelon Hulsebos, Yiming Lin, J. D. Zamfirescu-Pereira, Harrison Chase, Will Fu-Hinthorn, Aditya G. Parameswaran, and Eugene Wu. 2024b. spade: Synthesizing Data Quality Assertions for Large Language Model Pipelines. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4173–4186. https://doi.org/10.14778/3685800.3685835
- Shankar et al. (2024c) Shreya Shankar, J.D. Zamfirescu-Pereira, Bjoern Hartmann, Aditya Parameswaran, and Ian Arawjo. 2024c. Who Validates the Validators? Aligning LLM-Assisted Evaluation of LLM Outputs with Human Preferences. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 131, 14 pages. https://doi.org/10.1145/3654777.3676450
- Shute et al. (2024) Jeff Shute, Shannon Bales, Matthew Brown, Jean-Daniel Browne, Brandon Dolphin, Romit Kudtarkar, Andrey Litvinov, Jingchi Ma, John Morcos, Michael Shen, David Wilhite, Xi Wu, and Lulan Yu. 2024. SQL Has Problems. We Can Fix Them: Pipe Syntax In SQL. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4051–4063. https://doi.org/10.14778/3685800.3685826
- Singhvi et al. (2024) Arnav Singhvi, Manish Shetty, Shangyin Tan, Christopher Potts, Koushik Sen, Matei Zaharia, and Omar Khattab. 2024. DSPy Assertions: Computational Constraints for Self-Refining Language Model Pipelines. arXiv:2312.13382 [cs.CL] https://arxiv.org/abs/2312.13382
- Stoica et al. (2024) Ion Stoica, Matei Zaharia, Joseph Gonzalez, Ken Goldberg, Koushik Sen, Hao Zhang, Anastasios Angelopoulos, Shishir G. Patil, Lingjiao Chen, Wei-Lin Chiang, and Jared Q. Davis. 2024. Specifications: The missing link to making the development of LLM systems an engineering discipline. arXiv:2412.05299 [cs.SE] https://arxiv.org/abs/2412.05299
- Stonebraker (1975) Michael Stonebraker. 1975. Implementation of integrity constraints and views by query modification. In Proceedings of the 1975 ACM SIGMOD International Conference on Management of Data (San Jose, California) (SIGMOD ’75). Association for Computing Machinery, New York, NY, USA, 65–78. https://doi.org/10.1145/500080.500091
- Tang et al. (2024) Liyan Tang, Philippe Laban, and Greg Durrett. 2024. MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 8818–8847. https://doi.org/10.18653/v1/2024.emnlp-main.499
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 24824–24837. https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
- Willard and Louf (2023) Brandon T. Willard and Rémi Louf. 2023. Efficient Guided Generation for Large Language Models. arXiv:2307.09702 [cs.CL] https://arxiv.org/abs/2307.09702
- Zha et al. (2023) Yuheng Zha, Yichi Yang, Ruichen Li, and Zhiting Hu. 2023. AlignScore: Evaluating Factual Consistency with A Unified Alignment Function. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 11328–11348. https://doi.org/10.18653/v1/2023.acl-long.634
- Zheng et al. (2024) Lianmin Zheng, Liangsheng Yin, Zhiqiang Xie, Chuyue Sun, Jeff Huang, Cody Hao Yu, Shiyi Cao, Christos Kozyrakis, Ion Stoica, Joseph E. Gonzalez, Clark Barrett, and Ying Sheng. 2024. SGLang: Efficient Execution of Structured Language Model Programs. arXiv:2312.07104 [cs.AI] https://arxiv.org/abs/2312.07104

### v2

@@
 Abstract.
-The emergence of AI-augmented Data Processing Systems (DPSs) has introduced powerful semantic operators that extend traditional data management capabilities with LLM-based processing.
-However, these systems face fundamental reliability (a.k.a.
-trust) challenges, as LLMs can generate erroneous outputs, limiting their adoption in critical domains.
-Existing approaches to LLM constraints—ranging from user-defined functions to constrained decoding—are fragmented, imperative, and lack semantics-aware integration into query execution.
-To address this gap, we introduce Semantic Integrity Constraints (SICs), a novel declarative abstraction that extends traditional database integrity constraints to govern and optimize semantic operators within DPSs.
-SICs integrate seamlessly into the relational model, allowing users to specify common classes of constraints (e.g., grounding and soundness) while enabling query-aware enforcement and optimization strategies.
-In this paper, we present the core design of SICs, describe their formal integration into query execution, and detail our conception of grounding constraints, a key SIC class that ensures factual consistency of generated outputs.
-In addition, we explore novel enforcement mechanisms, combining proactive (constrained decoding) and reactive (validation and recovery) techniques to optimize efficiency and reliability.
-Our work establishes SICs as a foundational framework for trustworthy, high-performance AI-augmented data processing, paving the way for future research in constraint-driven optimizations, adaptive enforcement, and enterprise-scale deployments.
+AI-augmented data processing systems (DPSs) integrate large language models (LLMs) into query pipelines, allowing powerful semantic operations on structured and unstructured data.
+However, the reliability (a.k.a.
+trust) of these systems is fundamentally challenged by the potential for LLMs to produce errors, limiting their adoption in critical domains.
+To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries.
+SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both proactive and reactive enforcement strategies.
+We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems.
+Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs.
+To guide and evaluate the vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
 1. Introduction
 Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data.
-In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Shankar et al., 2024a; Liu et al., 2025; Anderson et al., 2025; Lu et al., 2025).
+In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025).
 We broadly refer to these systems as AI-augmented data processing systems (DPSs).
@@
 These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance.
-For these systems to be fully adopted, guardrails must be placed around semantic operators.
-LLM constraints—also known as assertions or guardrails—have been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023).
+Without guardrails, it will be difficult to fully adopt these systems.
+LLM constraints—also known as assertions or guardrails—have recently been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023; TruLens, 2021; Zheng et al., 2023).
 These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively.
-A related area of work involves constrained decoding, where constraints are applied to the LLM during its generation phase to ensure the output matches a specified structure (Beurer-Kellner et al., 2023; Dong et al., 2024b; guidance, 2022; OpenAI, 2024b; Willard and Louf, 2023; Zheng et al., 2024).
-While these existing approaches have proven their utility in enhancing the reliability of LLM outputs, they are imperative, fragmented, and lack awareness of operator semantics, limiting their usability and efficiency.
-To address these limitations, we introduce Semantic Integrity Constraints (SICs) as a unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs.
+A related approach is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023).
+While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
+We propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs.
 Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators.
-Similar to traditional integrity constraints, SICs make the specification of constraints for predominant use cases easy and more advanced cases possible.
-They are declarative, allowing for a rich set of enforcement and optimization techniques, such as constraint pushdown, constrained decoding, fused evaluation, and adaptive retries.
-In addition, SICs are expressed in the same language that defines data processing and integrate seamlessly into the relational model.
-These design decisions are in line with previous efforts that argue for statement and solution specifications on LLM-based system components (Stoica et al., 2024) as well as an (extended) relational model in AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Madden et al., 2024; Lu et al., 2025).
-The design of the right interface for SICs and the optimization of their enforcement present many exciting research opportunities and challenges.
-Different declarative interfaces have been proposed for semantic queries, with the majority adopting the relational model (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025).
-Although the community has not yet converged on an exact query language, we believe that the language for declaring semantic queries and SICs must be developed in tandem to maintain consistency and tight integration.
-Moreover, the constraint interface must enable users to declare SICs for the majority of use cases in a simple manner that does not require a deep understanding of the underlying implementation details.
-Towards this goal, the community will need to conduct user surveys to ensure that design decisions are grounded in real application requirements and do not introduce leaky abstractions.
-Lastly, optimizing SIC enforcement will require integrating modern techniques in LLM assertions and constrained decoding with traditional database query optimization methods.
-To meet ever-growing performance demands, novel approaches and extensions to existing solutions must also be developed.
-These solutions have the potential to accelerate the adoption of AI-augmented solutions in enterprise applications.
-This paper presents our vision for SICs and sketches a general SIC framework for relational data processing with unstructured data.
-Section 2 introduces our proposed constraint interface, followed by Section 3, a categorization of SICs that addresses the most common use cases for LLM constraints.
-In Section 4, we discuss the enforcement of SICs, detail various optimizations, and highlight multiple promising research directions.
-Section 5 focuses on grounding constraints, a class of SICs that we are implementing in VectraFlow (Lu et al., 2025), an AI-augmented data flow system developed at Brown University.
-Finally, we outline several interesting open problems in Section 6.
-2. Constraint Interface
-We propose an interface that adopts the relational model and enables users to define SICs with the same declarative language used for semantic queries.
-This approach aligns with traditional methods for expressing integrity constraints, where users declare the logic of a constraint without detailing its physical implementation, allowing the DPS to choose the most efficient way to enforce SICs.
-The majority of AI-augmented DPSs provide data flow interfaces, where users chain operators to define a query (Liu et al., 2025; Patel et al., 2024; Anderson et al., 2025).
-These systems offer a pandas-like interface in Python that exposes semantic operators separately from standard operators.
-This logical separation prevents declaring both semantic and standard predicates or functions within the same operator.
-However, recent work has introduced piped data flow syntax to SQL, enhancing its flexibility and extensibility (Shute et al., 2024).
-This new SQL dialect provides a basis for a unified interface between semantic and traditional queries, while also offering integrity constraints as first class citizens.
-Even though we design this interface for relational AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025), we expect that it can extend to DPSs with other data models too (Shankar et al., 2024a; Anderson et al., 2025).
-LABEL:lst:query shows an example query that incorporates semantic functionality into pipe SQL syntax.
-We first describe the high-level query and then discuss the SIC-specific code later in this and the following sections.
+Similar to their traditional counterparts, SICs make defining constraints for predominant use cases easy and more advanced cases possible, aligning with recent work that argues for specifications on LLM-based system components (Stoica et al., 2024).
+| Success Criterion | Description |
+| Expressiveness | SICs support a range of constraint types—including domain, grounding, and soundness—through a declarative interface (Section 3) |
+| Executable Semantics | Each constraint has precise execution semantics—including what is checked, when it is enforced, and what actions are taken on failure—ensuring semantically consistent system behavior and predictable outcomes for users (Section 4) |
+| System Integration | SICs integrate with query interfaces, optimizers, and execution engines in AI-augmented DPSs (Section 2) |
+| Performance Awareness | Enforcement strategy is selected based on cost and reliability across proactive or reactive methods (Section 4.2) |
+| Enterprise-Scale Applicability | SICs support observability, reuse, and auditability across workflows, with mechanisms for conflict detection and policy-driven enforcement at organizational scale (Sections 5 and 6) |
+This paper presents our vision for SICs and sketches a general framework for integrating them into AI-augmented DPSs.
+To evaluate the feasibility and impact of this vision, we define a set of design goals and success criteria: (1) expressiveness, to capture a broad range of common constraint types; (2) executable semantics, with well-defined runtime behavior; (3) system integration, via seamless support within data flow and query processing pipelines; (4) performance-aware enforcement, balancing reliability and cost; and (5) enterprise-scale applicability, including observability, constraint reuse, conflict detection, and auditability.
+These criteria, outlined in Table 1, were defined through our partnership with the Rhode Island Hospital.
+They guide our design and are reflected in the structure of the paper.
+Section 3.1 introduces the SIC interface, while Section 3.2 categorizes SICs across key enforcement classes aligned with real-world LLM use cases.
+Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms.
+Section 5 discusses observability components.
+Finally, Section 6 presents open challenges in large-scale constraint management.
+2. System Model
+We begin with an overview of the system architecture (Figure 1) for supporting SICs and defer a more detailed discussion of the core components to later sections.
+First, the user expresses their semantic query along with SICs via the query interface (Section 3.1).
+This interface forwards the query’s text to the query parser, which produces a corresponding logical query plan.
+Next, the logical optimizer (Section 4.1) applies logical rewrite rules to the initial plan, yielding an optimized logical plan.
+This plan is then passed to the physical optimizer (Section 4.2), which considers different physical implementations of the plan and selects the optimal one.
+The resulting optimized physical plan is then executed by the execution engine.
+To enable performance monitoring, an observability stack is integrated with the execution engine (Section 5).
+Finally, the constraint store recommends and analyzes SICs (Section 6).
+System architecture for SICs.
+Orange circles and turquoise diamonds represent semantic operators and SICs, respectively.
+Green squares indicate annotations that describe the physical implementation of each node.
+3. Constraint Specification
+We suggest defining SICs declaratively, allowing users to focus on constraint logic while the optimizer determines the best execution method.
+We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing.
+User surveys should be conducted to ensure usability.
+3.1. Interface
+To provide a unified interface for queries and SICs, users declare constraints in the query language.
+Specifically, we present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs.
+Though, SICs are compatible with any declarative AI-augmented DPS.
+The interface allows users to specify two types of information to the DPS for each constraint: (1) what the constraint is and (2) how the DPS should react when the constraint is violated.
+Declaring Constraints.
+SICs specify predicates either on generated tuple attributes from semantic operators or on the semantic operators themselves.
+We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case.
+The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT <attribute>, followed by a keyword that refers to the constraint class (Section 3.2) and any necessary parameters.
+A constraint on an operator is similar: ASSERT <operator>.
+Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
+Retry Thresholds.
+Users can define a retry threshold for each operator.
+When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached.
+To specify a retry threshold, users write RETRY <threshold> after the operator.
+Failure Modes.
+Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with operators and specify what the DPS should do when a failing output reaches the retry threshold.
+With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream.
+IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window).
+Lastly, ABORT aborts the entire query.
+To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
+End-To-End Example.
+LABEL:lst:query shows an example semantic query with various SICs.
 The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs).
-Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr, biological sex, and age.
-The query first canonicalizes the sex attribute’s value into “Male” or “Female” (line 2) because the value could have been in another representation, e.g., “M” or “F”.
-The SET expression is defined using a string prefixed with the letter “p”.
-This expression declares a prompt string literal (p-string), which specifies functions or predicates that the DPS evaluates with an LLM.
-Similar to Python’s formatted string literals (f-strings), p-strings can refer to other attributes in the query within curly braces.
-Lines 5 to 7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history.
-Because the EXTEND operator effectively invokes a semantic map to project new attributes, each projected attribute declares a name and type using the AS keyword.
-Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 9 to 10).
-Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the data flow (line 15).
-Further note that the order of operations specified in the query does not determine the execution order.
-For example, the DPS should execute the filter on line 9 before summarizing the patient’s medical history on line 8 to avoid excessive computation.
+Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr and dob.
+The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format.
+The expression is defined using a string prefixed with the letter “p”, which declares a prompt string.
+Prompt strings specify natural language functions or predicates over attributes and are evaluated using LLMs.
+Line 4 computes the patient’s age using standard SQL functions.
+Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history.
+Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 10-11).
+Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 16).
+The first constraint is declared on line 3, and the remaining constraints are specified on line 12 and onward.
+If the summarization step for med_hist_sum (line 8) fails its constraints (lines 13-15), the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraints (line 9).
+Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used.
+We explain each constraint in more detail in the next section.
 Semantic query with pipe SQL syntax and SICs.
-Crucially, our constraint interface extends the ASSERT operator provided by pipe SQL syntax, allowing users to declare SICs anywhere in the data flow.
-In the example query, assert operators are used in lines 4 and 11 and onward.
-These operators specify predicates that can refer to all attributes and operators defined upstream.
-This design enables the assert operator to validate values produced by attribute-generating semantic operators (e.g., maps, aggregates) as well as the presence or absence of tuples from nonattribute-generating operators (e.g., filters, joins, top-).
-Since the predicates of the declared assert operators may not be satisfied, the DPS can retry semantic operators in the case of failure.
-For each semantic operator, the user can specify a retry threshold, which is the maximum number of times the DPS can retry an operator with the same input.
-The user can also set a retry threshold that applies to all semantic operators in the data flow by default.
-In LABEL:lst:query, the retry threshold for the SET operator is defined in line 3, and all other operators use the default threshold.
-The user can also specify a failure mode (CONTINUE, DROP, or ABORT) for each assert operator, which determines how the DPS handles the failed input tuple(s).
-Similar to the retry threshold, the user can set a default failure mode.
-In the example query, line 4 defines the CONTINUE failure mode on the assert operator, while all other assert operators use the default.
-If a tuple does not satisfy the assert operator’s predicate, then the CONTINUE failure mode continues to propagate it downstream, while DROP drops it.
-Note that CONTINUE and DROP correspond to the (soft) Suggest and (hard) Assert constructs in DSPy Assertions (Singhvi et al., 2024), respectively.
-Lastly, ABORT aborts the entire query if a tuple does not satisfy the predicate.
-While the semantics of aborting are straightforward in a transactional context, in a data flow context, this means that the system will not only drop the tuple that violated the SIC but also discard all of its dependent tuples, such as those that would be in the same aggregation window.
-In any case, the DPS will log the appropriate errors and warnings when a tuple fails a SIC or when tuples are dropped to notify users of failures.
-3. Constraint Classes
-We categorize SICs into different classes (Table 1) that are designed to capture the predominant use cases for LLM constraints:
+3.2. Constraint Classes
+We categorize SICs into different classes (Table 2) that are designed to capture the predominant use cases for LLM constraints.
+We motivate each constraint class using concrete examples from the EHR query (among others) and briefly outline their implementation.
+Domain Constraints.
+In many queries, users want the generated attributes to belong to a specified domain.
+This includes use cases such as enforcing that a value is of a certain data type or structure, from a set of values, or less than a certain length.
+For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values.
+We extend traditional domain constraints to semantic data processing.
+In LABEL:lst:query, lines 3 and 15 specify domain constraints, restricting the dob to the regex representation of YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively.
+The type definitions (e.g., STRING) of generated attributes implicitly define domain constraints.
+This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
+Inclusion and Exclusion (IE) Constraints.
+Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles.
+For example, given a semantic map that summarizes an EHR, users may want to exclude all patient PII from the summary while including the names of attending physicians.
+IE constraints address this class of predicates.
+In the example query, line 14 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary.
+When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match.
+Users can also specify IE constraints using strings, regular expressions, or even a collection of them (e.g., a set of toxic words to exclude from the output).
+In such cases, deterministic methods are used to evaluate matches.
+Grounding Constraints.
+Semantic operators, such as maps or aggregates, can extract data or summarize from their inputs.
+However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data.
+Grounding constraints enable users to enforce such invariants on generated attributes.
+For example, line 12 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR.
+Line 13 asserts that the summarized med_hist_sum is factually consistent with the source EHR.
+In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr.
+Formally, if a grounding constraint is defined on an attribute, the DPS will recursively enforce grounding constraints on all ancestor attributes in its lineage.
+Thus, each constraint only needs to compare its operator’s input and output to ensure grounding in the source data.
+A grounding constraint is implemented in one of two ways, depending on whether the specified attribute is generated via an extractive (map) or abstractive (map or aggregate) operator.
+In the extractive case, the constraint is implemented deterministically and ensures that the operator’s output is precisely contained in its input.
+In the abstractive case, the constraint is implemented stochastically (e.g., with an LLM or a fact-checking model (Tang et al., 2024)) and ensures that the operator’s summarized output is factually consistent with the input.
+Users can annotate relevant operators with the EXTRACTIVE (lines 5-7) or ABSTRACTIVE (line 8) keywords to inform the DPS of their semantics; otherwise, the DPS infers these annotations by reasoning about the operator’s prompt string with an LLM.
 | Class | Description | Examples |
 | Domain | Value is from a specified domain | Data types, structured output, multiple choices, numerical ranges, or length constraints |
-| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHRs |
-| Grounding | Value is grounded in the source values | Extracted results or summaries are factual w.r.t. original EHRs |
+| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHR summaries |
+| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. original EHRs |
 | Soundness | LLM’s reasoning process is logically sound | Predictions of patient diagnoses are based on sound reasoning |
@@
 | Assertions | Arbitrary predicate holds in the data flow | All included URLs in generated research reports lead to valid websites |
-Domain Constraints.
-In the context of SICs, domain constraints remain similar to their traditional counterpart, checking whether an attribute’s value is from a specified domain.
-For example, a domain constraint can enforce that a value is of a certain data type or structure, within a numerical range, from a finite set of possible values, or less than a certain length.
-(Type definitions of generated attributes from semantic operators implicitly specify domain constraints.) In the example query from LABEL:lst:query, the first conjunct in line 4 and the conjunct in line 14 explicitly define domain constraints, restricting the sex to either “Male” or “Female” and the summary’s length to less than 1000 characters, respectively.
-This class of SICs comprise the majority of constraint use cases (Liu et al., 2024).
-Importantly, the DPS can enforce domain constraints programmatically without relying on an LLM validator, providing correctness guarantees on their evaluation.
-Inclusion and Exclusion (IE) Constraints.
-IE constraints determine if particular terms, content, topics, or abstract features like tones and styles are included or excluded from an attribute’s value.
-For instance, these constraints ensure that an attending physician’s name is present in an EHR, while avoiding the inclusion of PII.
-Users can specify these constraints with a set of values that indicate what must be included or excluded.
-This set can be derived from a fixed collection or another attribute’s value.
-The user can opt for checks based on exact keyword matches or semantic similarity, even using regular expressions for matches on emails or phone numbers.
-Users can also define IE constraints with a p-string for content, topics, tones, or styles, as shown in line 13 of the example query, which excludes test results from the medical history summary.
-In this case, the DPS employs an LLM to enforce the constraint, while lightweight models can be used to handle PII or toxic content checks, such as NER models, LSTMs, or feature-based classifiers.
-Grounding Constraints.
-These constraints target both extractive and abstractive use cases, where users want to check whether the output values of semantic operators are grounded in the source values.
-For instance, line 11 from the example query ensures that the extracted physical exam and lab results appear verbatim in the original EHR (extractive use case).
-Similarly, line 12 validates that the generated summary of the patient’s medical history is factually consistent with the EHR (abstractive use case).
-We discuss grounding constraints in greater detail in Section 5.
 Soundness Constraints.
-These constraints verify whether the underlying LLM’s reasoning process is logically sound in that its premises are true and its reasoning steps are valid.
-This class of constraints targets nonattribute-generating semantic operators and semantic transformations from one or more values to a new value.
-For such cases, it is nontrivial to simply evaluate predicates on the operator’s output without invoking an LLM validator to redo the task.
-Instead of this similarly error-prone approach, the DPS uses an LLM validator to check the underlying LLM’s reasoning process, e.g., chain-of-thought (CoT) (Wei et al., 2022).
-For example, consider the semantic filter (defined on lines 9 and 10), which selects tuples where the patient is predicted to have sepsis based on their medical condition.
+To check whether the reasoning of operators themselves is logically sound, users can declare soundness constraints.
+Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-.
+To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning process by modifying the operator to output its reasoning steps, e.g., chain-of-thought (CoT) (Wei et al., 2022).
+Line 16 declares a soundness constraint on the sepsis filter.
 Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity).
-This retrieved information is considered to outline sepsis symptoms and the diagnostic rationale.
+This retrieved information is considered to describe the symptoms of sepsis and the diagnostic rationale.
 Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result.
@@
 The premises are followed by the reasoning steps that lead to the selection result.
-Then, the input and output of the filter’s LLM are given to an LLM validator to ensure that the premises are grounded in the input and that the reasoning steps are valid.
-This step is defined with the assert operator in line 15.
-Similarly, a soundness constraint is specified in the second conjunct of line 4, which ensures that sex is canonicalized correctly.
-If the underlying LLM already performs well in a closed-book setting and the retrieved documents do not provide enough context to produce a valid answer, another approach is to have the LLM generate its CoT first and then verify each proposed premise with the input tuple(s) or documents retrieved afterwards.
-The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
-This is an avenue for future research.
+Then, the input and output of the filter’s LLM are given to an LLM validator (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid.
+The DPS can also use more sophisticated CoT prompting and verification techniques, which use structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
 Relevance Constraints.
-These constraints validate if an attribute’s value is relevant for the given semantic operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user.
+These constraints validate if an attribute’s value is relevant for the given operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user.
 Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default.
-We assume that this is the case in the example query; e.g., the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than their social history.
-Relevance constraints are evaluated by passing the operator’s prompt, input, and output to an LLM validator for verification.
+We assume that this is the case in the example query.
+For example, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history.
+These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.
 Assertions.
@@
 4. Constraint Enforcement
-Once users define SICs, the DPS will identify the most efficient enforcement method.
-The DPS initially moves all specified SICs to the earliest evaluation point to minimize unnecessary computation, akin to predicate pushdown in logical query optimization.
-If a SIC involves checking attributes, the DPS will move the constraint to the earliest data flow point where all attributes are defined, usually right after a semantic map or aggregate operator, since these are the only attribute-generating operators in our extended relational model.
-For SICs that verify operator processing logic, like soundness constraints, the DPS positions them at their associated semantic operator.
-Additionally, the DPS converts specified SICs into text and appends them to the LLM’s prompt, ensuring that the semantic operator’s LLM is aware of the constraints on its results.
-At this point, the DPS will follow either a proactive or reactive constraint enforcement strategy.
-4.1. Proactive Enforcement
-The DPS can proactively enforce certain types of SICs by pushing their enforcement into the LLM’s decoding process via constrained decoding (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022).
-In doing so, the output of the LLM will be guaranteed to satisfy the specified constraints by construction; thus, requiring no further checks on its output.
-Provided the constrained decoding overhead remains low, the DPS should proactively enforce SICs where feasible to optimize performance.
-Concerns around distorting the LLM’s distribution can be addressed with recent methods as well (Park et al., 2024).
-Grammar-Based Constrained Decoding.
-An effective method for using constrained decoding to ensure specific SICs is by encoding them as a formal grammar.
-This is similar to how libraries (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) allow users to provide a JSON Schema (Org, 2010) to which the LLM’s output adheres.
-These libraries convert the JSON Schema into a context-free grammar or regular expression and apply a push-down or finite automaton, respectively, to guide the output.
-This allows the DPS to enforce certain domain constraints.
-Beyond exploring more efficient grammar-based decoding methods, open challenges include determining which SICs can be expressed as grammars and how to represent and compile multiple SICs into a grammar if they are defined on the same attribute.
-Direct Constrained Decoding.
-An alternative constrained decoding method is to enforce SICs directly, without translating them into a formal grammar.
-Although we do not anticipate that this will allow the DPS to enforce all SICs, we expect it to enable a broader range of constraints than grammar-based methods.
-Previous research (Beurer-Kellner et al., 2023) introduced a direct, more flexible approach, but is significantly slower than current grammar-based techniques (Zheng et al., 2024).
-Whether this more flexible enforcement method can be implemented to efficiently enforce SICs is an open problem.
-4.2. Reactive Enforcement and Recovery Actions
-The DPS can also reactively enforce SICs by first detecting a constraint violation and then performing a recovery action.
-4.2.1. Violation Detection
-To detect a SIC violation, the DPS simply evaluates its associated predicate and checks if the predicate returns false.
-However, if multiple SICs are specified in the data flow, there are many opportunities to optimize their collective evaluation.
+Constraint enforcement consists of two phases: detection and recovery.
+First, a constraint detects any violations in the operator’s output.
+Upon detection of a violation, the operator attempts to recover.
+This distinction aids in reasoning about optimizing constraint enforcement, as discussed in this section.
+4.1. Logical Optimization
+The logical optimizer takes as input a logical query plan represented as a data flow graph, where each node is either an operator or a constraint.
+The goal of the logical optimizer is to minimize the cost of the plan by applying rewrite rules to semantic operators and constraints.
+Here, we refer to cost as any arbitrary cost metric, such as latency or monetary cost.
+To keep the presentation focused, we only discuss logical optimizations for constraints.
+Constraint Pushdown.
+First, the optimizer moves all constraints to their earliest evaluation point in the data flow (i.e., immediately following their corresponding operators).
+This allows the DPS to detect and recover from constraint violations as soon as possible rather than unnecessarily processing erroneous outputs.
+The optimizer also appends the textual representation of each constraint to their corresponding operator’s LLM prompt, ensuring that the underlying model is aware of the constraints on its results.
 Constraint Reordering.
-If the DPS needs to execute multiple SICs sequentially (e.g., if the user specifies the constraints on the same attribute), then it will reorder the constraints to reduce execution costs.
-This is similar to predicate reordering in traditional database optimization.
-For example, the DPS can reorder SICs based on the cost and selectivity of each SIC’s predicate (Abadi et al., 2003).
-The cost of a SIC can be based on a single metric, such as execution time, or even a linear combination of cost metrics.
-The selectivity of a SIC is the probability that an output tuple fails the associated predicate.
-The DPS can learn the cost and selectivity of each SIC via offline or online methods by computing lightweight statistics.
-Constraint Fusion.
-The DPS will further improve performance by merging SICs together into a single execution unit for certain scenarios.
-Akin to operator fusion, the DPS will combine code-based SICs together to reduce intermediate data materialization and unnecessary memory traffic.
-The DPS will also do the same for LLM-based SICs, where several constraints are evaluated by a single LLM invocation.
-Though, as more LLM-based constraints are fused together, the accuracy of results from the LLM validator is likely to degrade due to the increase in task complexity and input length (Levy et al., 2024).
-As such, finding the optimal number of constraints to combine that maintains target accuracy metrics while minimizing execution costs is an open problem.
-4.2.2. Violation Recovery
-To recover from a SIC violation, the DPS will retry the semantic operator with the failed input tuple(s) by using prompt or model-based techniques to improve the output quality.
-We outline a few approaches for retrying semantic operators.
-These approaches can be combined in different ways, and it is interesting to explore whether the DPS can learn their optimal combination with respect to different target metrics, such as accuracy, latency, and monetary cost.
-Few-Shot Prompting.
-A common approach for retrying a semantic operator is to append the erroneous output and corresponding feedback message to the previous prompt on each failure (Singhvi et al., 2024; Shankar et al., 2024a).
-The feedback message can be a static error message specified for a code-based predicate or part of the response from an LLM validator.
-This refinement process provides the underlying LLM negative few-shot examples of the given task for in-context learning (Brown et al., 2020).
-A complementary approach is to use the top- most similar positive few-shot examples, which are input-output pairs that previously satisfied all the specified SICs for the semantic operator.
-Importantly, examples are retrieved on the basis of their similarity to the LLM’s current input.
-The DPS can use a cache with a vector index to efficiently store and retrieve these examples.
-For each retry, the number of retrieved examples can increase to improve the quality of the LLM’s output.
-We believe that this approach will be particularly useful in data flow systems, where queries are mostly static.
-There are several challenges in our setting, e.g., determining ’s initial value and increment size for each retry, and example cache management policies.
-Chain-of-Thought Prompting.
-When retrying a semantic operator, the DPS can also prompt the underlying LLM to use more reasoning steps in its CoT.
-By effectively scaling up inference-time computation, recent work has shown that long CoT in reasoning models generally improves the accuracy of responses (OpenAI, 2024a; DeepSeek-AI, 2025).
-For each retry, the DPS can thus prompt the LLM to use increasingly longer CoT to generate better outputs.
-It is unclear how long the initial CoT should be and by how much to increase it on each retry.
-Model-Based Improvements.
-To improve the results of semantic operators, the DPS can invoke a more powerful LLM to generate a more accurate output, or use multiple LLMs to adopt a majority voting approach, e.g., where the same input is sent to increasingly more LLMs on each retry.
-In simple cases, such as when the operator outputs a classification label or a numerical value, the DPS can programmatically consider the outputs of each LLM and return the most common result or their average, respectively.
-In more complex scenarios, such as when the operator generates a summary, the DPS can use a separate LLM to consider all the outputs and produce the final result.
-An open avenue for exploration is to compare the accuracy of these two methods when various cost metrics are fixed.
-5. Grounding Constraints
-In this section, we describe grounding constraints in more detail to illustrate the design decisions required to integrate SICs.
-These constraints are being implemented in VectraFlow (Lu et al., 2025), an AI-augmented data flow system that we are developing at Brown.
-However, we aim to keep the presentation general since the ideas proposed here are applicable to other AI-augmented DPSs as well.
-We first discuss preliminaries related to grounding constraints and then describe their implementation in the rest of the section.
-To understand the rationale for grounding constraints, we note a key distinction: an LLM’s response comes from its internal (parametric) knowledge, external (nonparametric) knowledge, or a combination of both.
-In an AI-augmented DPS, this external knowledge includes input tuples to the system or tuples retrieved from stored tables, termed source tuples, with their elements called source values.
-These values, considered factual, are essential to grounding constraints.
-Across extractive and abstractive contexts, users seek to confirm if a value generated by a semantic operator is anchored in source values.
-Grounding constraints facilitate these verifications.
-5.1. Enforcement
-When a grounding constraint is specified for an attribute, the DPS recursively applies checks to all ancestor attributes in its lineage.
-These grounding checks are necessary because the attribute may have been derived through a series of intermediate semantic operators from the source value(s).
-The DPS evaluates these checks immediately after attributes are generated, ensuring that the operator’s output is grounded in its input.
-In the example query from LABEL:lst:query, med_hist_sum is derived from med_hist, which, in turn, is derived from ehr.
-Since the grounding constraint specified on line 12 for med_hist_sum recursively defines a grounding check on med_hist, the query does not need to explicitly specify a grounding constraint on med_hist, though doing so is equivalent.
-The notion of grounding differs depending on whether the operator is extractive or abstractive.
-For extractive operators, grounded output is precisely contained within the input.
-For abstractive operators, grounded output is factually consistent with the input.
-The DPS cannot inherently identify whether an operator is extractive or abstractive; hence, users must annotate extractive semantic maps.
-Without such annotations, attribute-generating operators in the data flow are presumed abstractive if they produce an attribute within the grounding constraint’s lineage.
-In the example query, we see the EXTRACTIVE keyword included for the extractive operations on lines 5 to 7, while no annotation is present for the abstractive operation that summarizes the medical history on line 8.
-Evaluating grounding checks on extractive and abstractive semantic operators necessitates different check types.
-The DPS will automatically assess the appropriate checks based on each operator’s semantics.
-We will now discuss these grounding check types.
-5.2. Types of Grounding Checks
-We propose three types of grounding checks with different verification mechanisms, targeting either extractive or abstractive tasks.
-Exact Match.
-A grounding check based on exact match verifies whether an output attribute’s value is exactly contained within the input value.
-These checks are designed for extractive semantic map operations, where users require exact guarantees on output values, such as phys_exam, lab_res, and med_hist in the example query.
-The DPS can further optimize exact match verification by building inverted indices over stored relations.
-However, in data flow systems, evaluating these checks on values derived from input data streams are not amenable to such optimizations.
-Factual Consistency Score.
-These checks evaluate whether an output value is grounded in the input value(s) by computing factual consistency scores with a specialized model.
-Each sentence in an output value is separately given with the input value(s) to the factual consistency model, and the resulting score determines whether the sentence is grounded.
-While these models are stochastic, they can validate summaries from abstractive semantic maps (e.g., med_hist_sum) and aggregates.
-Models like MiniCheck have also been shown to be more efficient than their LLM counterparts and reach GPT-4 accuracy (Tang et al., 2024).
-LLM Validator.
-LLM-based grounding checks also target abstractive use cases, invoking an LLM validator to determine grounding.
-The LLM validator receives the output value and input value(s), and it returns a factual consistency score or a boolean value.
-Additionally, the LLM validator provides a feedback message, which can help the semantic operator output more grounded responses when retried.
-In fact, the DPS can combine the LLM-based check with the score-based check, where it invokes the LLM validator only if the factual consistency model classifies the output as ungrounded.
-6. Open Problems
-In this section, we present additional open problems that pave the way for promising research opportunities.
-Adaptive Enforcement.
-A DPS can strategically manage performance and risk by dynamically applying SICs.
-By tracking SIC violations for each semantic operator, it determines when checks are necessary, thus reducing overhead.
-Violations incur a (normalized) penalty score reflecting their risk level, and these scores are combined into a sum.
-If this sum exceeds a set threshold, a retry is initiated to maintain critical integrity.
-This approach allows the system to adapt on the basis of each operator’s reliability, concentrating on critical validations while minimizing disruptions when operator confidence is high.
-Another approach involves adaptively enforcing SICs based on target metrics (e.g., accuracy, F1 score).
-LLM Input Verification.
-While our discussion of SICs has mainly emphasized validating semantic operator outputs, it is equally crucial to validate the inputs, which may come from external sources (including a stored table).
-This step is necessary to avoid security risks, such as (i) prompt attacks (AI, 2024; Llama, 2024; Microsoft, 2024; Rebedea et al., 2023) or (ii) prevent unauthorized access to dark data (IBM, 2023), which refers to unprotected data hidden in a vector database that should not be used because it can pose security, privacy, or compliance risks due to its sensitive contents.
-An open challenge here is developing robust and scalable mechanisms to perform real-time, context-aware input verification that can effectively identify and filter malicious or unauthorized data before it reaches the LLM.
-Enterprise-Wide Automated Constraint Verification.
-To enable a “check all” feature for semantic queries, each semantic operator in the query path must be automatically verified.
-Strategically determining what checks to use, where and how to place them involves examining data flow semantics and lineage to pinpoint where operator outputs affect subsequent results.
-The types of checks, which range from basic range checks to LLM validators, should match each operator’s semantics and risk of its output.
-Although our discussion has focused on single queries, enterprise settings often involve numerous queries that share common domain-specific constraints.
-Therefore, an important challenge is developing a mechanism to automatically capture and manage these recurring constraints in a constraint store and apply them consistently across all relevant semantic queries, ensuring enterprise-wide data integrity.
-To further enhance this process, the DPS should log all SIC violations, coupled with user feedback on incorrect validation results.
-This approach facilitates continuous LLM refinement through training, fine-tuning, or in-context learning, ultimately bolstering the accuracy of enterprise-wide constraint verification.
-Leveraging LLM Progress.
-As LLM reliability and efficiency advance, they will influence SIC design and related trade-offs.
-Increases in LLM accuracy allow for more adaptivity in SIC enforcement.
-A critical question is: How do we determine when LLM accuracies warrant this shift?
-Additionally, as LLMs become more efficient, using advanced ensemble methods (e.g., confidence-based aggregation) to further improve reliability as well as maintaining LLM diversity to prevent shared errors become important.
+The optimizer then reorders the SICs for the same operator.
+Following prior work (Abadi et al., 2003), the reordering is based on the cost and selectivity of each constraint, where selectivity is defined as the probability that a tuple satisfies the constraint’s predicate.
+Cost and selectivity estimates can be computed from sampled data.
+We discuss more details on estimation in Section 4.2.2.
+4.2. Physical Optimization
+Given an optimized logical plan, the physical optimizer considers various implementations of the plan and returns the optimal one.
+We begin with a discussion of different approaches to constraint enforcement, and then outline ideas for optimized plan selection.
+4.2.1. Enforcement Implementations
+Constraints can be enforced in multiple ways, typically categorized as reactive or proactive, based on whether enforcement occurs after or during response generation, respectively.
+While the DPS can enforce all constraints reactively, only certain ones can be enforced proactively.
+Reactive Enforcement.
+After the LLM produces its entire response, there are different methods to detect a constraint violation.
+For example, a grounding constraint on an extractive operator can simply check if the operator’s output attribute is contained in the input.
+On the other hand, a grounding constraint on an abstractive operator can invoke a single LLM judge (Zheng et al., 2023), a fact-checking model (Tang et al., 2024), a compound AI system (Chern et al., 2023), or anything in between.
+Similarly, there are different approaches for an operator to recover from a violation by retrying.
+The common approach used in current systems is iterative refinement with feedback (Singhvi et al., 2024; Shankar et al., 2024a; Madaan et al., 2023).
+Furthermore, the operator can retrieve additional semantically similar few-shot examples (Brown et al., 2020) of prior input-output pairs that satisfied the constraint.
+These examples can be stored in a cache built on a vector index.
+Other approaches include retrying with more powerful models, larger ensemble sizes, or higher thinking budgets (Anthropic, 2025; Google, 2025).
+Proactive Enforcement.
+Proactively enforcing constraints can lower costs significantly.
+For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, identifying and masking invalid tokens to ensure compliance by construction.
+Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token validation.
+However, constrained decoding is not limited to grammar-based approaches.
+For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints.
+Yet, open source libraries currently do not offer this functionality.
+An open challenge is to identify which other SICs can be enforced via constrained decoding.
+Moreover, developing a framework to efficiently apply various constraints on different output attributes also remains unresolved.
+Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024).
+Nonetheless, it provides a foundation for future exploration.
+Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions.
+This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
+Besides deterministic constrained decoding, the DPS can proactively enforce stochastic constraints by asynchronously detecting violations in partial streamed outputs.
+This approach is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint.
+For instance, the DPS can proactively enforce a grounding constraint on an abstractive operator by asynchronously validating generated summary sentences with a fact-checking model (Tang et al., 2024).
+Upon a violation, the operator can backtrack to the decoding step right before the violation.
+Alternative techniques like re-prompting with partial feedback are also worth considering to address this problem.
+4.2.2. Optimized Plan Selection
+For a fixed set of constraints and their implementations, the physical optimizer can adapt existing techniques (Liu et al., 2025b; Russo et al., 2025) to select an optimal query plan defined by user-specified cost or reliability (a.k.a.
+quality) thresholds (omitted from the query interface for brevity).
+Specifically, rather than measuring the reliability of an operator’s implementation in terms of validation data or the best implementation from prior beliefs (e.g., a “champion model”), reliability is instead defined by the proportion of output tuples that satisfy all of the operator’s constraints (after necessary retries).
+However, the challenge is selecting each constraint’s implementation to begin with.
+As discussed in Section 4.2.1, there are many ways to detect a constraint violation, each with different cost-reliability trade-offs.
+Formally, the cost of a constraint’s implementation is the overhead of detecting a violation in a single output.
+The reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation.
+A deterministic constraint is completely reliable (i.e., 100% precision and recall), while a stochastic constraint is not completely reliable, due to invoking models such as LLMs.
+To choose the best constraint implementation, we must estimate costs and reliability for each option.
+While cost estimation is often simple and can be done during runtime with sampled data, it might not always be necessary.
+For instance, proactively enforcing a deterministic constraint through constrained decoding is usually quicker than doing it reactively.
+Assessing the reliability of stochastic constraints is more challenging.
+Unlike deterministic constraints that are always reliable, the reliability of stochastic constraints depends on (1) how they are implemented, (2) the output tuples from upstream operators in the query plan, and (3) the input tuples to the query.
+Moreover, obtaining robust reliability estimates for constraint implementations can be difficult, as they often need ground-truth labels from human annotators.
+We present several strategies to address this issue.
+One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime.
+This approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates.
+Since metrics are averaged across diverse query contexts, they do not effectively reflect interactions with upstream operators.
+Furthermore, data and workload drift may make these estimates obsolete.
+It is also generally costly for organizations to annotate execution traces.
+An alternative approach is to develop a shared foundation model for organizations, where given representations of (1) a constraint’s implementation, (2) the implementation of the query plan up until the constraint, and (3) the underlying data, the model predicts the precision and recall for the constraint’s implementation.
+Organizations can optionally fine-tune the model for their use cases.
+Recent work on foundation database models (Wehrstein et al., 2025) are applicable here, with extensions to address the complexities of semantic processing.
+A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., expected confidence) of a constraint’s implementation at runtime based on sampled data.
+Rather than defining precision and recall thresholds, users specify confidence thresholds instead.
+Although many conventional classifiers inherently provide confidence scores, they can be derived from LLMs by applying softmax to the “True” and “False” token logits.
+The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability.
+Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks.
+Though, further research is needed to obtain highly-calibrated constraints.
+5. Observability
+After a user submits a query to the DPS, the observability stack enables the user to understand the performance of their query’s operators and constraints.
+The stack has three components.
+The observability interface allows the user to view relevant metrics (e.g., cost, reliability) and inspect query results.
+The observability store maintains metrics and outputs from each stage in the query.
+Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint.
+While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread in both academia and industry (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), their reliability remains uncertain.
+As a result, labeling interfaces remain essential.
+A key open challenge lies in ensuring that the observability store can support the proposed interfaces efficiently at scale and that these interfaces are effectively designed.
+We detail each component of the stack below.
+Observability Interface.
+Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity.
+However, to observe the reliability for stochastic constraints, the user must submit a labeling request for their query.
+Once the labeling is complete, the interface displays the precision and recall of each constraint.
+To avoid confusion, the interface indicates whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well.
+The observability interface also allows the user to inspect individual output tuples.
+A tuple is annotated if any upstream operator violated a constraint but CONTINUEed to propagate the error.
+The user can view the tuple’s lineage for more fine-grained debugging.
+Observability Store.
+The observability store is modeled as additional relations in the DPS.
+It maintains the lineage of each query, enabling systematic debugging and performance monitoring.
+To obtain human annotations, the DPS can draw concepts from crowd-sourced databases (Franklin et al., 2011; Marcus et al., 2011; Park et al., 2012), such as crowd-sourced attributes and user interface generation; however, the DPS need not implement these ideas in their full generality.
+For example, suppose the DPS stores a relation where each tuple represents an invocation of a constraint.
+The relation includes attributes for the constraint’s input and output (i.e., predicted label) among others (e.g., query ID).
+In particular, the relation also has a crowd-sourced attribute for the true label (initially null).
+When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample of the relation) the precision and recall of each constraint for the given query ID.
+The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
+Labeling Interface.
+A key aspect of crowd-sourced databases is their ability to automatically create user interfaces for crowd workers based on declarative queries.
+Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs.
+The annotator’s role is to determine if the inputs meet the constraint.
+To avoid bias, the original predicted label should be hidden.
+6. Enterprise-Wide Constraints
+The introduction of SICs into an enterprise environment presents both new opportunities and challenges.
+To effectively manage SICs within an organization, we propose the development of a constraint store and outline its key desired features.
+Auto-Guardrails and Recommendations.
+Although declaring SICs is relatively straightforward, determining the applicable constraints is more difficult.
+This requires anticipating the potential failure cases for each semantic operator in the query.
+Recent systems have been developed to assist users in formulating constraints (Shankar et al., 2024b, c), but these approaches typically generate suggestions based solely on the contents of the user’s query.
+In enterprise settings, constraint recommendations can be greatly enriched.
+Organizations typically enforce domain-specific rules that necessitate certain constraints for all pertinent queries.
+For instance, a hospital might require excluding PII in queries unless essential for analysis.
+Moreover, different teams within the same organization might use similar queries with shared constraints.
+Thus, the constraint store needs to automatically capture and manage common constraints organization-wide, facilitating their reuse and recommendation in related queries.
+We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries.
+This strategy boosts the efficiency of similarity searches (Malkov and Yashunin, 2020) for identifying relevant constraints, which can then be refined by an LLM to match the user’s context.
+Conflicts.
+Conflicts can arise when queries contain multiple SICs (e.g., from accumulating shared constraints across an organization’s workloads).
+For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy both constraints without redaction or abstraction.
+Likewise, domain constraints over disjoint value sets can lead to inevitable failures.
+The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings.
+Users can address conflicts by adjusting priorities (e.g., making one constraint as soft), relaxing scope, or merging constraints as needed.
+This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
 7. Related Work
 Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975).
-Since then, constraints have been supported by all major databases to ensure that modifications maintain data consistency.
-SICs build on top of traditional integrity constraints, improving the reliability of the semantic operators themselves by protecting against erroneous outputs from the underlying LLMs.
-In ML, assertions were introduced as an abstraction to monitor and improve models, invoking UDFs over model outputs to flag possible errors (Kang et al., 2020).
-Recent works adopt similar ideas by further leveraging ML models and LLM validators to define guardrails on LLMs (Rebedea et al., 2023; Dong et al., 2024a; AI, 2023a, b; Index, 2023), with DSPy Assertions introducing an iterative refinement approach to improve output quality (Singhvi et al., 2024).
-Methods for validating an LLM’s CoT have been developed as well (Ling et al., 2023).
-In addition to these reactive constraint enforcement approaches, a related area of work uses constrained decoding to proactively enforce constraints (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022; Beurer-Kellner et al., 2023).
-SICs consolidate these various approaches into a unified and declarative abstraction designed for AI-augmented DPSs.
-Prior works like spade (Shankar et al., 2024b), which automatically synthesizes assertions, and EvalGen (Shankar et al., 2024c), which aligns LLM validators with human preferences, are complimentary to SICs and can be integrated to enhance usability and reliability.
-Several recent systems have augmented traditional database operators with the semantic capabilities of LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025; Lu et al., 2025; Shankar et al., 2024a).
-So far, DocETL (Shankar et al., 2024a) is the only AI-augmented DPS to provide validation as a first class citizen.
+Since then, constraints have been supported by all major databases.
+ML assertions were introduced to monitor and enhance model performance by applying UDFs to model outputs to find potential errors (Kang et al., 2020).
+Building upon this idea, recent developments broaden assertions for LLMs by employing LLMs as judges and ML classifiers for both offline evaluation (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024).
+Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts).
+These tools complement SICs and can suggest declarative constraints to boost enforcement reliability and efficiency.
+Our vision for enterprise constraints also draws inspiration from these efforts.
+DSPy Assertions enable users to define assertions in LLM pipelines using UDFs, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024).
+While SICs use iterative refinement, they do not currently include prompt optimization capabilities, although they can be added.
+To the best of our knowledge, no framework provides a fully declarative method for optimized constraint enforcement.
+A related area of work is constrained decoding, which proactively enforces constraints so that the LLM’s response adheres to them by construction (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023).
+Current frameworks focus on grammar-based approaches; however, similar to earlier work (Beurer-Kellner et al., 2023), we view constrained decoding as a more general approach that is not limited to expressing constraints as grammars.
+Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025).
+So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a).
 They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs.
-We further their approach by proposing a declarative, relational interface for the easy specification of common constraint classes and their optimized enforcement.
-Many methods have been developed to evaluate the factual consistency of LLM responses with respect to provided source documents.
-This problem is also known as document-level textual entailment.
-In addition to an exact keyword match approach that is only appropriate for extractive use cases, recent methods for evaluating factual consistency in abstractive tasks are largely categorized as those that use LLM validators and those that use specialized factual consistency models.
-The latter models are specifically trained to evaluate factual consistency, primarily in the context of summarization (Tang et al., 2024; Zha et al., 2023; Kryscinski et al., 2020; Goyal and Durrett, 2020; Laban et al., 2022).
-These models provide factual consistency scores and are more efficient than LLM-based methods, with some even reaching GPT-4 accuracy (Tang et al., 2024; Zha et al., 2023).
-Regardless of the specific method for evaluating factual consistency, the interface provided by SICs is designed to be agnostic of the underlying implementation.
+However, their approach to validation is largely imperative.
+Palimpzest does not support constraints, but they allow users to specify quality thresholds to select optimized physical plans (Liu et al., 2025b; Russo et al., 2025).
 8. Conclusions
-SICs are a unified abstraction that extend traditional database integrity constraints to enhance the reliability of AI-augmented DPSs.
-They provide a declarative, relational interface to easily specify common classes of constraint while maintaining the flexibility to support more sophisticated use cases.
-Crucially, this declarative interface enables the DPS to apply a wide range of highly optimized implementations.
-By adopting this principled approach, SICs open new, grounded research avenues towards the goal of leveraging AI-augmented DPSs for critical workloads where reliability is paramount.
+SICs offer a unified, declarative foundation for improving the reliability of AI-augmented data processing.
+By generalizing traditional integrity constraints to the semantic behaviors of LLM-augmented operators, SICs bridge the gap between declarative query processing and the uncertainty of generative models.
+Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse.
+Through this framework, we outline a principled path toward building auditable, trustworthy AI-augmented systems capable of supporting mission-critical workflows.
+SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.
 Acknowledgements.
@@
 - AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
-- AI (2024) Guardrails AI. 2024. Detect Jailbreak. https://hub.guardrailsai.com/validator/guardrails/detect_jailbreak
 - AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
+- Aiken et al.
+(1992) Alexander Aiken, Jennifer Widom, and Joseph M Hellerstein.
+1992.
+Behavior of database production rules: Termination, confluence, and observable determinism.
+In Proceedings of the 1992 ACM SIGMOD international conference on Management of data.
+59–68.
 - Anderson et al.
@@
 In CIDR.
+- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. https://www.anthropic.com/news/visible-extended-thinking
 - Astrahan et al.
@@
 Transactions on Machine Learning Research (2023).
+- Chern et al.
+(2023) I-Chun Chern, Steffi Chern, Shiqi Chen, Weizhe Yuan, Kehua Feng, Chunting Zhou, Junxian He, Graham Neubig, and Pengfei Liu.
+2023.
+FacTool: Factuality Detection in Generative AI–A Tool Augmented Framework for Multi-Task and Multi-Domain Scenarios.
+arXiv:2307.13528 [cs.CL] https://arxiv.org/abs/2307.13528
 - Codd (1970) E. F. Codd. 1970. A relational model of data for large shared data banks. Commun. ACM 13, 6 (June 1970), 377–387. https://doi.org/10.1145/362384.362685
-- DeepSeek-AI (2025) DeepSeek-AI. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948 [cs.CL] https://arxiv.org/abs/2501.12948
+- Dai et al.
+(2024) Hanjun Dai, Bethany Yixin Wang, Xingchen Wan, Bo Dai, Sherry Yang, Azade Nova, Pengcheng Yin, Phitchaya Mangpo Phothilimthana, Charles Sutton, and Dale Schuurmans.
+2024.
+UQE: A Query Engine for Unstructured Databases.
+In Advances in Neural Information Processing Systems, A.
+Globerson, L.
+Mackey, D.
+Belgrave, A.
+Fan, U.
+Paquet, J.
+Tomczak, and C.
+Zhang (Eds.), Vol.
+37.
+Curran Associates, Inc., 29807–29838.
+https://proceedings.neurips.cc/paper_files/paper/2024/file/34b3a40ec9752c1ae48fe85fef8fe8dc-Paper-Conference.pdf
+- Dayal et al.
+(1988) Umeshwar Dayal, Barbara Blaustein, Alex Buchmann, Upen Chakravarthy, Meichun Hsu, R Ledin, Dennis McCarthy, Arnon Rosenthal, Sunil Sarin, Michael J.
+Carey, et al.
+1988.
+The HiPAC project: Combining active databases and timing constraints.
+ACM Sigmod Record 17, 1 (1988), 51–70.
 - Dong et al.
@@
 https://doi.org/10.1145/1282480.1282484
-- Goyal and Durrett (2020) Tanya Goyal and Greg Durrett.
-2020.
-Evaluating Factuality in Generation with Dependency-level Entailment.
-In Findings of the Association for Computational Linguistics: EMNLP 2020, Trevor Cohn, Yulan He, and Yang Liu (Eds.).
-Association for Computational Linguistics, Online, 3592–3603.
-https://doi.org/10.18653/v1/2020.findings-emnlp.322
+- Franklin et al.
+(2011) Michael J.
+Franklin, Donald Kossmann, Tim Kraska, Sukriti Ramesh, and Reynold Xin.
+2011.
+CrowdDB: answering queries with crowdsourcing.
+In Proceedings of the 2011 ACM SIGMOD International Conference on Management of Data (Athens, Greece) (SIGMOD ’11).
+Association for Computing Machinery, New York, NY, USA, 61–72.
+https://doi.org/10.1145/1989323.1989331
+- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
 - guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
+- Guo et al.
+(2017) Chuan Guo, Geoff Pleiss, Yu Sun, and Kilian Q.
+Weinberger.
+2017.
+On calibration of modern neural networks.
+In Proceedings of the 34th International Conference on Machine Learning - Volume 70 (Sydney, NSW, Australia) (ICML’17).
+JMLR.org, 1321–1330.
+- Hanson (1996) Eric N. Hanson. 1996. The design and implementation of the Ariel active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 1 (1996), 157–172.
 - Held et al.
@@
 https://doi.org/10.1145/1499949.1500029
-- IBM (2023) IBM. 2023. What is dark data? Retrieved 2025-02-27 from https://www.ibm.com/think/topics/dark-data
 - Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
+- Kadavath et al.
+(2022) Saurav Kadavath, Tom Conerly, Amanda Askell, Tom Henighan, Dawn Drain, Ethan Perez, Nicholas Schiefer, Zac Hatfield-Dodds, Nova DasSarma, Eli Tran-Johnson, Scott Johnston, Sheer El-Showk, Andy Jones, Nelson Elhage, Tristan Hume, Anna Chen, Yuntao Bai, Sam Bowman, Stanislav Fort, Deep Ganguli, Danny Hernandez, Josh Jacobson, Jackson Kernion, Shauna Kravec, Liane Lovitt, Kamal Ndousse, Catherine Olsson, Sam Ringer, Dario Amodei, Tom Brown, Jack Clark, Nicholas Joseph, Ben Mann, Sam McCandlish, Chris Olah, and Jared Kaplan.
+2022.
+Language Models (Mostly) Know What They Know.
+arXiv:2207.05221 [cs.CL] https://arxiv.org/abs/2207.05221
 - Kang et al.
@@
 481–496.
-- Kryscinski et al.
-(2020) Wojciech Kryscinski, Bryan McCann, Caiming Xiong, and Richard Socher.
-2020.
-Evaluating the Factual Consistency of Abstractive Text Summarization.
-In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.).
-Association for Computational Linguistics, Online, 9332–9346.
-https://doi.org/10.18653/v1/2020.emnlp-main.750
-- Laban et al.
-(2022) Philippe Laban, Tobias Schnabel, Paul N.
-Bennett, and Marti A.
-Hearst.
-2022.
-SummaC: Re-Visiting NLI-based Models for Inconsistency Detection in Summarization.
-Transactions of the Association for Computational Linguistics 10 (2022), 163–177.
-https://doi.org/10.1162/tacl_a_00453
-- Levy et al.
-(2024) Mosh Levy, Alon Jacoby, and Yoav Goldberg.
+- Kipf and Welling (2017) Thomas N.
+Kipf and Max Welling.
+2017.
+Semi-Supervised Classification with Graph Convolutional Networks.
+In International Conference on Learning Representations.
+https://openreview.net/forum?id=SJU4ayYgl
+- Koo et al.
+(2024) Terry Koo, Frederick Liu, and Luheng He.
 2024.
-Same Task, More Tokens: the Impact of Input Length on the Reasoning Performance of Large Language Models.
-In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.).
-Association for Computational Linguistics, Bangkok, Thailand, 15339–15353.
-https://doi.org/10.18653/v1/2024.acl-long.818
+Automata-based constraints for language model decoding.
+In First Conference on Language Modeling.
+https://openreview.net/forum?id=BDBdblmyzY
+- Kuchnik et al.
+(2023) Michael Kuchnik, Virginia Smith, and George Amvrosiadis.
+2023.
+Validating Large Language Models with ReLM.
+In Proceedings of Machine Learning and Systems, D.
+Song, M.
+Carbin, and T.
+Chen (Eds.), Vol.
+5.
+Curan, 457–476.
+https://proceedings.mlsys.org/paper_files/paper/2023/file/93c7d9da61ccb2a60ac047e92787c3ef-Paper-mlsys2023.pdf
 - Ling et al.
@@
 - Liu et al.
-(2025) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano.
-2025.
+(2025b) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano.
+2025b.
 Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing.
 In CIDR.
+- Liu et al.
+(2025a) Hongfu Liu, Hengguan Huang, Xiangming Gu, Hao Wang, and Ye Wang.
+2025a.
+On Calibration of LLM-based Guard Models for Reliable Content Moderation.
+In The Thirteenth International Conference on Learning Representations.
+https://openreview.net/forum?id=wUbum0nd9N
 - Liu et al.
@@
 https://doi.org/10.1145/3613905.3650756
-- Llama (2024) Meta Llama. 2024. Prompt Guard. https://www.llama.com/docs/model-cards-and-prompt-formats/prompt-guard/
 - Lu et al. (2025) Duo Lu, Siming Feng, Jonathan Zhou, Franco Solleza, Malte Schwarzkopf, and Uğur Çetintemel. 2025. VectraFlow: Integrating Vectors into Stream Processing. In CIDR.
@@
 arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
-- Madden et al.
-(2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska.
-2024.
-Databases Unbound: Querying All of the World’s Bytes with AI.
+- Madaan et al.
+(2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, Shashank Gupta, Bodhisattwa Prasad Majumder, Katherine Hermann, Sean Welleck, Amir Yazdanbakhsh, and Peter Clark.
+2023.
+SELF-REFINE: iterative refinement with self-feedback.
+In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23).
+Curran Associates Inc., Red Hook, NY, USA, Article 2019, 61 pages.
+- Malkov and Yashunin (2020) Yu A.
+Malkov and D.
+A.
+Yashunin.
+2020.
+Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs.
+IEEE Transactions on Pattern Analysis and Machine Intelligence 42, 4 (2020), 824–836.
+https://doi.org/10.1109/TPAMI.2018.2889473
+- Marcus et al. (2011) Adam Marcus, Eugene Wu, David R. Karger, Samuel Madden, and Robert C. Miller. 2011. Crowdsourced Databases: Query Processing with People. In CIDR.
+- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. https://openai.com/index/introducing-text-and-code-embeddings/
+- OpenAI (2024) OpenAI. 2024. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
+- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
+- Park et al.
+(2012) Hyunjung Park, Hector Garcia-Molina, Richard Pang, Neoklis Polyzotis, Aditya Parameswaran, and Jennifer Widom.
+2012.
+Deco: a system for declarative crowdsourcing.
 Proc.
 VLDB Endow.
-17, 12 (Aug.
-2024), 4546–4554.
-https://doi.org/10.14778/3685800.3685916
-- Microsoft (2024) Microsoft. 2024. Prompt Shields. https://learn.microsoft.com/en-us/azure/ai-services/content-safety/concepts/jailbreak-detection
-- OpenAI (2024a) OpenAI. 2024a. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/
-- OpenAI (2024b) OpenAI. 2024b. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
-- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
-- Park et al.
-(2024) Kanghee Park, Jiayu Wang, Taylor Berg-Kirkpatrick, Nadia Polikarpova, and Loris D'Antoni.
-2024.
-Grammar-Aligned Decoding.
-In Advances in Neural Information Processing Systems, A.
-Globerson, L.
-Mackey, D.
-Belgrave, A.
-Fan, U.
-Paquet, J.
-Tomczak, and C.
-Zhang (Eds.), Vol.
-37.
-Curran Associates, Inc., 24547–24568.
-https://proceedings.neurips.cc/paper_files/paper/2024/file/2bdc2267c3d7d01523e2e17ac0a754f3-Paper-Conference.pdf
+5, 12 (Aug.
+2012), 1990–1993.
+https://doi.org/10.14778/2367502.2367555
 - Patel et al.
@@
 https://doi.org/10.18653/v1/2023.emnlp-demo.40
+- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. arXiv:1908.10084 [cs.CL] https://arxiv.org/abs/1908.10084
+- Russo et al.
+(2025) Matthew Russo, Sivaprasad Sudhir, Gerardo Vitagliano, Chunwei Liu, Tim Kraska, Samuel Madden, and Michael Cafarella.
+2025.
+Abacus: A Cost-Based Optimizer for Semantic Operator Systems.
+arXiv:2505.14661 [cs.DB] https://arxiv.org/abs/2505.14661
+- Santhanam et al.
+(2024) Keshav Santhanam, Deepti Raghavan, Muhammad Shahir Rahman, Thejas Venkatesh, Neha Kunjal, Pratiksha Thaker, Philip Levis, and Matei Zaharia.
+2024.
+ALTO: An Efficient Network Orchestrator for Compound AI Systems.
+In Proceedings of the 4th Workshop on Machine Learning and Systems (Athens, Greece) (EuroMLSys ’24).
+Association for Computing Machinery, New York, NY, USA, 117–125.
+https://doi.org/10.1145/3642970.3655844
 - Shankar et al.
@@
 https://doi.org/10.18653/v1/2024.emnlp-main.499
+- TruLens (2021) TruLens. 2021. TruLens. https://www.trulens.org/
+- Urban and Binnig (2024) Matthias Urban and Carsten Binnig. 2024. CAESURA: Language Models as Multi-Model Query Planners. In CIDR.
+- Veličković et al.
+(2018) Petar Veličković, Guillem Cucurull, Arantxa Casanova, Adriana Romero, Pietro Liò, and Yoshua Bengio.
+2018.
+Graph Attention Networks.
+In International Conference on Learning Representations.
+https://openreview.net/forum?id=rJXMpikCZ
+- Wang and Li (2025) Jiayi Wang and Guoliang Li. 2025. AOP: Automated and Interactive LLM Pipeline Orchestration for Answering Complex Queries. In CIDR.
+- Wehrstein et al. (2025) Johannes Wehrstein, Carsten Binnig, Fatma Özcan, Shobha Vasudevan, Yu Gan, and Yawen Wang. 2025. Towards Foundation Database Models. In CIDR.
 - Wei et al.
-(2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou.
+(2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H.
+Chi, Quoc V.
+Le, and Denny Zhou.
 2022.
 Chain-of-thought prompting elicits reasoning in large language models.
-In Advances in Neural Information Processing Systems, S.
-Koyejo, S.
-Mohamed, A.
-Agarwal, D.
-Belgrave, K.
-Cho, and A.
-Oh (Eds.), Vol.
-35.
-Curran Associates, Inc., 24824–24837.
-https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
+In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22).
+Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
+- Widom (1996) Jennifer Widom. 1996. The starburst active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 4 (1996), 583–595.
 - Willard and Louf (2023) Brandon T. Willard and Rémi Louf. 2023. Efficient Guided Generation for Large Language Models. arXiv:2307.09702 [cs.CL] https://arxiv.org/abs/2307.09702
-- Zha et al.
-(2023) Yuheng Zha, Yichi Yang, Ruichen Li, and Zhiting Hu.
+- Xu et al.
+(2019) Keyulu Xu, Weihua Hu, Jure Leskovec, and Stefanie Jegelka.
+2019.
+How Powerful are Graph Neural Networks?.
+In International Conference on Learning Representations.
+https://openreview.net/forum?id=ryGs6iA5Km
+- Zheng et al.
+(2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric P.
+Xing, Hao Zhang, Joseph E.
+Gonzalez, and Ion Stoica.
 2023.
-AlignScore: Evaluating Factual Consistency with A Unified Alignment Function.
-In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.).
-Association for Computational Linguistics, Toronto, Canada, 11328–11348.
-https://doi.org/10.18653/v1/2023.acl-long.634
+Judging LLM-as-a-judge with MT-bench and Chatbot Arena.
+In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23).
+Curran Associates Inc., Red Hook, NY, USA, Article 2020, 29 pages.
 - Zheng et al.

### v3

@@
 To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries.
-SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both proactive and reactive enforcement strategies.
+SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both reactive and proactive enforcement strategies.
 We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems.
 Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs.
-To guide and evaluate the vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
+To guide and evaluate our vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
 1. Introduction
 Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data.
-In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025).
+In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025; Agarwal and Huang, 2025; Madden et al., 2024).
 We broadly refer to these systems as AI-augmented data processing systems (DPSs).
@@
 These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance.
-Without guardrails, it will be difficult to fully adopt these systems.
-LLM constraints—also known as assertions or guardrails—have recently been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023; TruLens, 2021; Zheng et al., 2023).
-These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively.
-A related approach is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023).
+Without guardrails, it will be difficult to adopt these systems fully.
+LLM assertions have actively been explored for offline evaluations (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024; Loula et al., 2025).
+These assertions invoke custom code, ML classifiers, or LLM judges to check an LLM’s output with respect to various criteria.
+A related technique is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023).
 While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
-We propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs.
+To address these limitations, we propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs.
 Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators.
@@
 | System Integration | SICs integrate with query interfaces, optimizers, and execution engines in AI-augmented DPSs (Section 2) |
-| Performance Awareness | Enforcement strategy is selected based on cost and reliability across proactive or reactive methods (Section 4.2) |
+| Performance Awareness | Enforcement strategy is selected based on cost and reliability across reactive and proactive methods (Section 4.2) |
 | Enterprise-Scale Applicability | SICs support observability, reuse, and auditability across workflows, with mechanisms for conflict detection and policy-driven enforcement at organizational scale (Sections 5 and 6) |
@@
 They guide our design and are reflected in the structure of the paper.
-Section 3.1 introduces the SIC interface, while Section 3.2 categorizes SICs across key enforcement classes aligned with real-world LLM use cases.
+Section 2 outlines the proposed system architecture.
+Section 3 introduces the SIC interface and categorizes SICs across key enforcement classes aligned with real-world LLM use cases.
 Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms.
@@
 Finally, Section 6 presents open challenges in large-scale constraint management.
-2. System Model
+2. System Architecture
 We begin with an overview of the system architecture (Figure 1) for supporting SICs and defer a more detailed discussion of the core components to later sections.
@@
 Finally, the constraint store recommends and analyzes SICs (Section 6).
-System architecture for SICs.
-Orange circles and turquoise diamonds represent semantic operators and SICs, respectively.
-Green squares indicate annotations that describe the physical implementation of each node.
 3. Constraint Specification
@@
 We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing.
-User surveys should be conducted to ensure usability.
+Additional surveys should be conducted to ensure usability.
 3.1. Interface
 To provide a unified interface for queries and SICs, users declare constraints in the query language.
-Specifically, we present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs.
-Though, SICs are compatible with any declarative AI-augmented DPS.
-The interface allows users to specify two types of information to the DPS for each constraint: (1) what the constraint is and (2) how the DPS should react when the constraint is violated.
+We present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs; however, SICs are compatible with any declarative AI-augmented DPS.
 Declaring Constraints.
@@
 We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case.
-The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT <attribute>, followed by a keyword that refers to the constraint class (Section 3.2) and any necessary parameters.
-A constraint on an operator is similar: ASSERT <operator>.
+The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT, followed by a predicate involving the attribute’s name, the type of constraint (Section 3.2), and any necessary parameters.
+A constraint on an operator is similar.
 Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
 Retry Thresholds.
-Users can define a retry threshold for each operator.
+Users can define a retry threshold for each constraint.
 When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached.
-To specify a retry threshold, users write RETRY <threshold> after the operator.
+To specify a retry threshold, users write RETRY <threshold> after the constraint.
 Failure Modes.
-Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with operators and specify what the DPS should do when a failing output reaches the retry threshold.
+Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with constraints and specify what the DPS should do when a failing output reaches the retry threshold.
 With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream.
 IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window).
-Lastly, ABORT aborts the entire query.
+Lastly, ABORT cancels the entire query.
 To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
@@
 The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format.
-The expression is defined using a string prefixed with the letter “p”, which declares a prompt string.
-Prompt strings specify natural language functions or predicates over attributes and are evaluated using LLMs.
+The expression involves a string prefixed with the letter “p”, which defines a prompt string.
+Prompt strings can specify natural language functions or predicates over attributes and are evaluated using LLMs.
 Line 4 computes the patient’s age using standard SQL functions.
 Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history.
-Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 10-11).
-Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 16).
-The first constraint is declared on line 3, and the remaining constraints are specified on line 12 and onward.
-If the summarization step for med_hist_sum (line 8) fails its constraints (lines 13-15), the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraints (line 9).
+Lastly, the query filters tuples by selecting patients likely to have sepsis (line 9).
+Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 14).
+The first constraint is declared on line 3, and the remaining constraints are specified on line 10 and onward.
+If the summarization step for med_hist_sum (line 8) fails the constraint on line 12, the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraint again (line 13).
 Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used.
@@
 Semantic query with pipe SQL syntax and SICs.
+System architecture for SICs.
+Orange circles and turquoise diamonds represent semantic operators and SICs, respectively.
+Green squares indicate annotations that describe the physical implementation of each node.
 3.2. Constraint Classes
@@
 Domain Constraints.
-In many queries, users want the generated attributes to belong to a specified domain.
-This includes use cases such as enforcing that a value is of a certain data type or structure, from a set of values, or less than a certain length.
+In many queries, users want generated attributes to belong to a specified domain.
+This includes use cases such as enforcing that a value conforms to a certain data type or structure, belongs to a set of values, or is less than a certain length.
 For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values.
 We extend traditional domain constraints to semantic data processing.
-In LABEL:lst:query, lines 3 and 15 specify domain constraints, restricting the dob to the regex representation of YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively.
-The type definitions (e.g., STRING) of generated attributes implicitly define domain constraints.
+In LABEL:lst:query, line 3 and the second conjunct on line 11 specify domain constraints, restricting the dob to the regex for YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively.
+Type declarations (e.g., STRING) of generated attributes implicitly define domain constraints.
 This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
@@
 Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles.
-For example, given a semantic map that summarizes an EHR, users may want to exclude all patient PII from the summary while including the names of attending physicians.
+For example, given a semantic map that summarizes an EHR, users may want to exclude all the patient’s PII from the summary while including the names of attending physicians.
 IE constraints address this class of predicates.
-In the example query, line 14 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary.
+Line 12 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary.
 When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match.
-Users can also specify IE constraints using strings, regular expressions, or even a collection of them (e.g., a set of toxic words to exclude from the output).
+Users can also specify IE constraints using strings, regexes, or even a collection of them (e.g., a set of toxic words to exclude from the output).
 In such cases, deterministic methods are used to evaluate matches.
@@
 However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data.
-Grounding constraints enable users to enforce such invariants on generated attributes.
-For example, line 12 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR.
-Line 13 asserts that the summarized med_hist_sum is factually consistent with the source EHR.
+Grounding constraints enforce such invariants.
+For example, line 10 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR, while the first conjunct on line 11 asserts that the summarized med_hist_sum is factually consistent with the source EHR.
 In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr.
@@
 | Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHR summaries |
-| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. original EHRs |
+| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. source EHRs |
 | Soundness | LLM’s reasoning process is logically sound | Predictions of patient diagnoses are based on sound reasoning |
@@
 Soundness Constraints.
-To check whether the reasoning of operators themselves is logically sound, users can declare soundness constraints.
+To check whether the reasoning of an operator is logically sound, users can declare soundness constraints.
 Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-.
-To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning process by modifying the operator to output its reasoning steps, e.g., chain-of-thought (CoT) (Wei et al., 2022).
-Line 16 declares a soundness constraint on the sepsis filter.
+To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning, e.g., chain-of-thought (CoT) (Wei et al., 2022).
+As an example, line 14 declares a soundness constraint on the sepsis filter.
 Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity).
@@
 The premises are followed by the reasoning steps that lead to the selection result.
-Then, the input and output of the filter’s LLM are given to an LLM validator (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid.
-The DPS can also use more sophisticated CoT prompting and verification techniques, which use structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
+Then, the input and output of the filter’s LLM are given to an LLM judge (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid.
+The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats (Ling et al., 2023).
 Relevance Constraints.
@@
 We assume that this is the case in the example query.
-For example, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history.
+For instance, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history.
 These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.
 Assertions.
-Like assertions in traditional integrity constraints, these constraints enable users to specify arbitrary predicates that must hold in the data flow.
-The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the primitives described previously.
+Like assertions in traditional integrity constraints, this class of constraints enables users to specify arbitrary predicates that must hold in the data flow.
+The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the classes described previously.
 4. Constraint Enforcement
@@
 First, a constraint detects any violations in the operator’s output.
-Upon detection of a violation, the operator attempts to recover.
-This distinction aids in reasoning about optimizing constraint enforcement, as discussed in this section.
+If a violation is detected, the operator attempts to recover.
+This distinction helps us reason about optimizing constraint enforcement, as discussed in this section.
 4.1. Logical Optimization
@@
 4.2. Physical Optimization
-Given an optimized logical plan, the physical optimizer considers various implementations of the plan and returns the optimal one.
-We begin with a discussion of different approaches to constraint enforcement, and then outline ideas for optimized plan selection.
+Given an optimized logical plan, the physical optimizer considers different implementations of the plan and returns the optimal one.
+We first discuss approaches to constraint enforcement and then outline ideas for optimized plan selection.
 4.2.1. Enforcement Implementations
-Constraints can be enforced in multiple ways, typically categorized as reactive or proactive, based on whether enforcement occurs after or during response generation, respectively.
-While the DPS can enforce all constraints reactively, only certain ones can be enforced proactively.
+Constraints are enforced reactively or proactively, based on whether enforcement occurs after or during response generation.
+While all constraints can be enforced reactively, only some can be enforced proactively.
 Reactive Enforcement.
@@
 Proactively enforcing constraints can lower costs significantly.
-For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, identifying and masking invalid tokens to ensure compliance by construction.
-Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token validation.
+For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, masking invalid tokens to ensure compliance by construction.
+Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token masking.
 However, constrained decoding is not limited to grammar-based approaches.
-For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints.
-Yet, open source libraries currently do not offer this functionality.
+For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints for extractive operators.
+Yet, we are not aware of any existing library that offers this functionality.
 An open challenge is to identify which other SICs can be enforced via constrained decoding.
-Moreover, developing a framework to efficiently apply various constraints on different output attributes also remains unresolved.
+Moreover, developing a framework to apply efficiently various constraints on different output attributes also remains unresolved.
 Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024).
-Nonetheless, it provides a foundation for future exploration.
+Nonetheless, it provides a foundation for future investigation.
 Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions.
 This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
-Besides deterministic constrained decoding, the DPS can proactively enforce stochastic constraints by asynchronously detecting violations in partial streamed outputs.
-This approach is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint.
-For instance, the DPS can proactively enforce a grounding constraint on an abstractive operator by asynchronously validating generated summary sentences with a fact-checking model (Tang et al., 2024).
-Upon a violation, the operator can backtrack to the decoding step right before the violation.
-Alternative techniques like re-prompting with partial feedback are also worth considering to address this problem.
+Besides deterministic constrained decoding, the DPS can proactively enforce some stochastic constraints on partial outputs.
+For instance, the DPS can enforce a grounding constraint on an abstractive operator by invoking a fact-checking model (Tang et al., 2024) after each sentence is generated.
+While recent research (Loula et al., 2025) explores these ideas, they do not focus on efficient optimizations, such as asynchronously detecting violations in partial streamed outputs.
+This optimization is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint.
+Upon a violation, the operator can backtrack to the decoding step right before the violation; though, alternative techniques like re-prompting with partial feedback are also worth considering.
 4.2.2. Optimized Plan Selection
@@
 quality) thresholds (omitted from the query interface for brevity).
-Specifically, rather than measuring the reliability of an operator’s implementation in terms of validation data or the best implementation from prior beliefs (e.g., a “champion model”), reliability is instead defined by the proportion of output tuples that satisfy all of the operator’s constraints (after necessary retries).
-However, the challenge is selecting each constraint’s implementation to begin with.
-As discussed in Section 4.2.1, there are many ways to detect a constraint violation, each with different cost-reliability trade-offs.
-Formally, the cost of a constraint’s implementation is the overhead of detecting a violation in a single output.
-The reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation.
-A deterministic constraint is completely reliable (i.e., 100% precision and recall), while a stochastic constraint is not completely reliable, due to invoking models such as LLMs.
-To choose the best constraint implementation, we must estimate costs and reliability for each option.
-While cost estimation is often simple and can be done during runtime with sampled data, it might not always be necessary.
-For instance, proactively enforcing a deterministic constraint through constrained decoding is usually quicker than doing it reactively.
-Assessing the reliability of stochastic constraints is more challenging.
-Unlike deterministic constraints that are always reliable, the reliability of stochastic constraints depends on (1) how they are implemented, (2) the output tuples from upstream operators in the query plan, and (3) the input tuples to the query.
-Moreover, obtaining robust reliability estimates for constraint implementations can be difficult, as they often need ground-truth labels from human annotators.
-We present several strategies to address this issue.
+Specifically, rather than measuring the reliability of an operator’s implementation in terms of an oracle, reliability is instead defined by adherence to all of the operator’s constraints (after necessary retries).
+The challenge is selecting each constraint’s implementation in the first place.
+As illustrated in Section 4.2.1, each constraint’s implementation has unique cost-reliability trade-offs, which allow users to specify additional thresholds to control them.
+While existing cost estimation approaches can be extended to account for each operator’s constraints, estimating the reliability of the constraints is a more difficult task.
+Formally, the reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation.
+A deterministic constraint is completely reliable (i.e., 100% precision and recall) and requires no estimation, while a stochastic constraint is not completely reliable due to invoking models.
+Estimating the reliability of stochastic constraint implementations is challenging.
+First, such estimates require ground-truth labels from human annotators.
+Furthermore, reliability is a function of both the implementation and its input, and this input, in turn, depends on the upstream query plan and the underlying data.
+We present some initial ideas for constraint reliability estimation.
 One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime.
-This approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates.
-Since metrics are averaged across diverse query contexts, they do not effectively reflect interactions with upstream operators.
+However, this approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates.
+Since metrics are aggregated across diverse query contexts, they do not effectively reflect interactions with upstream operators.
 Furthermore, data and workload drift may make these estimates obsolete.
-It is also generally costly for organizations to annotate execution traces.
+It is also costly for individual organizations to annotate execution traces.
 An alternative approach is to develop a shared foundation model for organizations, where given representations of (1) a constraint’s implementation, (2) the implementation of the query plan up until the constraint, and (3) the underlying data, the model predicts the precision and recall for the constraint’s implementation.
@@
 Recent work on foundation database models (Wehrstein et al., 2025) are applicable here, with extensions to address the complexities of semantic processing.
-A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., expected confidence) of a constraint’s implementation at runtime based on sampled data.
+A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., confidence) of a constraint’s implementation at runtime based on sampled data.
 Rather than defining precision and recall thresholds, users specify confidence thresholds instead.
-Although many conventional classifiers inherently provide confidence scores, they can be derived from LLMs by applying softmax to the “True” and “False” token logits.
+While many conventional classifiers inherently provide confidence scores, they can also be derived from LLMs by applying softmax to the “True” and “False” token logits.
 The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability.
 Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks.
-Though, further research is needed to obtain highly-calibrated constraints.
+Though, further research is needed to obtain highly-calibrated constraints across varying domains.
 5. Observability
@@
 Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint.
-While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread in both academia and industry (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), their reliability remains uncertain.
+While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), they are not entirely reliable.
 As a result, labeling interfaces remain essential.
-A key open challenge lies in ensuring that the observability store can support the proposed interfaces efficiently at scale and that these interfaces are effectively designed.
 We detail each component of the stack below.
@@
 Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity.
-However, to observe the reliability for stochastic constraints, the user must submit a labeling request for their query.
+However, to observe the reliability of stochastic constraints, the user must submit a labeling request for their query.
 Once the labeling is complete, the interface displays the precision and recall of each constraint.
-To avoid confusion, the interface indicates whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well.
-The observability interface also allows the user to inspect individual output tuples.
-A tuple is annotated if any upstream operator violated a constraint but CONTINUEed to propagate the error.
-The user can view the tuple’s lineage for more fine-grained debugging.
+The interface can also indicate whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well.
+Moreover, the observability interface allows the user to inspect individual output tuples.
+A tuple is flagged if any upstream operator violated a constraint but CONTINUEed to propagate the error.
+The user can also view the tuple’s lineage for more fine-grained debugging.
 Observability Store.
@@
 In particular, the relation also has a crowd-sourced attribute for the true label (initially null).
-When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample of the relation) the precision and recall of each constraint for the given query ID.
+When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample) the precision and recall of each constraint for the given query ID.
 The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
 Labeling Interface.
-A key aspect of crowd-sourced databases is their ability to automatically create user interfaces for crowd workers based on declarative queries.
+A key aspect of crowd-sourced databases is their ability to create user interfaces automatically for crowd workers based on declarative queries.
 Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs.
-The annotator’s role is to determine if the inputs meet the constraint.
-To avoid bias, the original predicted label should be hidden.
+The annotator’s role is to determine if the inputs satisfy the constraint.
 6. Enterprise-Wide Constraints
@@
 To effectively manage SICs within an organization, we propose the development of a constraint store and outline its key desired features.
-Auto-Guardrails and Recommendations.
-Although declaring SICs is relatively straightforward, determining the applicable constraints is more difficult.
+Recommendations.
+Although declaring SICs is relatively simple, determining the applicable constraints is more difficult.
 This requires anticipating the potential failure cases for each semantic operator in the query.
@@
 Moreover, different teams within the same organization might use similar queries with shared constraints.
-Thus, the constraint store needs to automatically capture and manage common constraints organization-wide, facilitating their reuse and recommendation in related queries.
+Thus, the constraint store needs to capture and manage common constraints automatically organization-wide, facilitating their reuse and recommendation in related queries.
 We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries.
-This strategy boosts the efficiency of similarity searches (Malkov and Yashunin, 2020) for identifying relevant constraints, which can then be refined by an LLM to match the user’s context.
+This approach enables similarity searches to identify relevant constraints, which can then be refined by a reranker and LLM to match the user’s context.
 Conflicts.
-Conflicts can arise when queries contain multiple SICs (e.g., from accumulating shared constraints across an organization’s workloads).
-For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy both constraints without redaction or abstraction.
+Conflicts can arise when queries contain multiple SICs, such as from accumulating shared constraints across an organization’s workloads.
+For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy the constraints without redaction or abstraction.
 Likewise, domain constraints over disjoint value sets can lead to inevitable failures.
 The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings.
-Users can address conflicts by adjusting priorities (e.g., making one constraint as soft), relaxing scope, or merging constraints as needed.
+Users can resolve conflicts by adjusting priorities (e.g., making one constraint as soft) or relaxing scope as needed.
 This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
@@
 Since then, constraints have been supported by all major databases.
-ML assertions were introduced to monitor and enhance model performance by applying UDFs to model outputs to find potential errors (Kang et al., 2020).
-Building upon this idea, recent developments broaden assertions for LLMs by employing LLMs as judges and ML classifiers for both offline evaluation (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024).
+ML assertions were introduced to monitor and enhance model performance by applying arbitrary functions to model outputs to find potential errors (Kang et al., 2020).
+Recent developments broaden assertions to LLMs by employing LLM judges and ML classifiers for offline evaluations (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024).
 Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts).
-These tools complement SICs and can suggest declarative constraints to boost enforcement reliability and efficiency.
+These tools complement SICs and can instead suggest declarative constraints to boost enforcement reliability and efficiency.
 Our vision for enterprise constraints also draws inspiration from these efforts.
-DSPy Assertions enable users to define assertions in LLM pipelines using UDFs, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024).
-While SICs use iterative refinement, they do not currently include prompt optimization capabilities, although they can be added.
-To the best of our knowledge, no framework provides a fully declarative method for optimized constraint enforcement.
-A related area of work is constrained decoding, which proactively enforces constraints so that the LLM’s response adheres to them by construction (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023).
-Current frameworks focus on grammar-based approaches; however, similar to earlier work (Beurer-Kellner et al., 2023), we view constrained decoding as a more general approach that is not limited to expressing constraints as grammars.
-Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025).
+DSPy Assertions enable users to implement assertions in LLM pipelines as functions, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024).
+While SICs use iterative refinement, they do not currently support prompt optimization capabilities, although they can be added.
+A related area of work is controlled generation, which encompasses constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) and model-based constraints (Loula et al., 2025).
+To the best of our knowledge, there is no unified framework that offers declarative and optimized constraint enforcement.
+Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025; Agarwal and Huang, 2025; Madden et al., 2024).
 So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a).
@@
 However, their approach to validation is largely imperative.
-Palimpzest does not support constraints, but they allow users to specify quality thresholds to select optimized physical plans (Liu et al., 2025b; Russo et al., 2025).
+LOTUS (Patel et al., 2024) and Palimpzest (Liu et al., 2025b; Russo et al., 2025) do not support constraints, but they allow users to specify quality thresholds relative to oracles.
 8. Conclusions
@@
 Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse.
-Through this framework, we outline a principled path toward building auditable, trustworthy AI-augmented systems capable of supporting mission-critical workflows.
+Through this framework, we outline a principled path toward building trustworthy AI-augmented systems capable of supporting mission-critical workflows.
 SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.
 Acknowledgements.
-We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback and contributions.References
+This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant Nos 2439559 and 2040433.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
+We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback.References
 - (1)
@@
 https://doi.org/10.1007/s00778-003-0095-z
-- AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
-- AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
+- Agarwal and Huang (2025) Arun Agarwal and Renee Huang.
+2025.
+Introducing Cortex AISQL: Reimagining SQL into AI Query Language for Multimodal Data.
+Retrieved 2025-07-08 from https://www.snowflake.com/en/blog/ai-sql-query-language/
+- AI (2023a) Guardrails AI. 2023a. Guardrails. Retrieved 2025-07-08 from https://github.com/guardrails-ai/guardrails
+- AI (2023b) LangChain AI. 2023b. LangChain. Retrieved 2025-07-08 from https://github.com/langchain-ai/langchain
 - Aiken et al.
@@
 In CIDR.
-- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. https://www.anthropic.com/news/visible-extended-thinking
+- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. Retrieved 2025-07-08 from https://www.anthropic.com/news/visible-extended-thinking
 - Astrahan et al.
@@
 https://doi.org/10.1147/rd.206.0560
-- Chen et al.
-(2023) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W.
-Cohen.
-2023.
-Program of Thoughts Prompting: Disentangling Computation from Reasoning for Numerical Reasoning Tasks.
-Transactions on Machine Learning Research (2023).
 - Chern et al.
@@
 https://doi.org/10.1145/1989323.1989331
-- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
-- guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
+- Google (2025) Google.
+2025.
+Gemini 2.5 brings enhanced reasoning to enterprise use cases.
+Retrieved 2025-07-08 from https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
+- guidance (2022) guidance. 2022. guidance-ai. Retrieved 2025-07-08 from https://github.com/guidance-ai/guidance
 - Guo et al.
@@
 https://doi.org/10.1145/1499949.1500029
-- Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
+- Index (2023) Llama Index. 2023. Llama Index. Retrieved 2025-07-08 from https://github.com/run-llama/llama_index
 - Kadavath et al.
@@
 https://doi.org/10.1145/3613905.3650756
+- Loula et al.
+(2025) João Loula, Benjamin LeBrun, Li Du, Ben Lipkin, Clemente Pasti, Gabriel Grand, Tianyu Liu, Yahya Emara, Marjorie Freedman, Jason Eisner, Ryan Cotterell, Vikash Mansinghka, Alexander K.
+Lew, Tim Vieira, and Timothy J.
+O’Donnell.
+2025.
+Syntactic and Semantic Control of Large Language Models via Sequential Monte Carlo.
+In The Thirteenth International Conference on Learning Representations.
+https://openreview.net/forum?id=xoXn62FzD0
 - Lu et al. (2025) Duo Lu, Siming Feng, Jonathan Zhou, Franco Solleza, Malte Schwarzkopf, and Uğur Çetintemel. 2025. VectraFlow: Integrating Vectors into Stream Processing. In CIDR.
-- Lyu et al.
-(2023) Qing Lyu, Shreya Havaldar, Adam Stein, Li Zhang, Delip Rao, Eric Wong, Marianna Apidianaki, and Chris Callison-Burch.
-2023.
-Faithful Chain-of-Thought Reasoning.
-arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
 - Madaan et al.
@@
 Curran Associates Inc., Red Hook, NY, USA, Article 2019, 61 pages.
-- Malkov and Yashunin (2020) Yu A.
-Malkov and D.
-A.
-Yashunin.
-2020.
-Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs.
-IEEE Transactions on Pattern Analysis and Machine Intelligence 42, 4 (2020), 824–836.
-https://doi.org/10.1109/TPAMI.2018.2889473
+- Madden et al.
+(2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska.
+2024.
+Databases Unbound: Querying All of the World’s Bytes with AI.
+Proc.
+VLDB Endow.
+17, 12 (Aug.
+2024), 4546–4554.
+https://doi.org/10.14778/3685800.3685916
 - Marcus et al. (2011) Adam Marcus, Eugene Wu, David R. Karger, Samuel Madden, and Robert C. Miller. 2011. Crowdsourced Databases: Query Processing with People. In CIDR.
-- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. https://openai.com/index/introducing-text-and-code-embeddings/
-- OpenAI (2024) OpenAI. 2024. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
-- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
+- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. Retrieved 2025-07-08 from https://openai.com/index/introducing-text-and-code-embeddings/
+- OpenAI (2024) OpenAI. 2024. Structured Outputs. Retrieved 2025-07-08 from https://platform.openai.com/docs/guides/structured-outputs/
+- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. Retrieved 2025-07-08 from https://github.com/json-schema-org/json-schema-spec
 - Park et al.
@@
 https://doi.org/10.18653/v1/2024.emnlp-main.499
-- TruLens (2021) TruLens. 2021. TruLens. https://www.trulens.org/
+- TruLens (2021) TruLens. 2021. TruLens. Retrieved 2025-07-08 from https://www.trulens.org/
 - Urban and Binnig (2024) Matthias Urban and Carsten Binnig. 2024. CAESURA: Language Models as Multi-Model Query Planners. In CIDR.